### v1

# Self-Employment as a Signal: Career Concerns with Hidden Firm Performance

Abstract
We analyze a dynamic labor market in which a worker with career concerns chooses each period between (i) self-employment that makes output publicly observable and (ii) employment at a firm that pays a flat wage but keeps individual performance hidden from outside observers. Output is binary and the worker is risk averse. The worker values future opportunities through a reputation for talent; firms may be benchmark (myopic) (ignoring the informational content of an application) or equilibrium (updating beliefs from the very act of applying). Three forces shape equilibrium: an insurance–information trade-off, selection by reputation, and inference from application decisions. We show that (i) an absorbing employment region exists in which low-reputation workers strictly prefer the firm’s insurance and optimally cease producing public information; (ii) sufficiently strong reputation triggers self-employment in order to generate public signals and preserve future outside options; and (iii) with equilibrium firms, application choices act as signals that shift hiring thresholds and wages even when in-firm performance remains opaque. Comparative statics deliver sharp, testable predictions for the prevalence of self-employment, the cyclicality of switching, and wage dynamics across markets with different degrees of performance transparency. The framework links classic career-concerns models to contemporary environments in which some tasks generate portable, public histories while firm tasks remain unobserved by the outside market (e.g., open-source contributions, freelancing platforms, or sales roles with standardized public metrics). Our results rationalize recent empirical findings on the value of public performance records and illuminate when opacity inside firms dampens or amplifies reputational incentives.
keywords:
career concerns , reputation , self-employment , dynamic signaling , adverse selection
JEL:
D82 , D83 , J24 , J31 , L26 , M51
journal: Working Paper
1 Introduction
Markets often see some but not all of what workers do. [^1] Think open-source commits and package releases for developers, contest entries or preprints for researchers, portfolio pieces for designers, client reviews for freelancers, and seller ratings on platforms. A developer’s open-source commits, a freelancer’s platform ratings, or a salesperson’s client-visible metrics are public; most day-to-day contributions inside firms are not. We study a dynamic environment in which a risk-averse worker cares about both current consumption and a reputation for talent that shapes future wages. Each period the worker chooses between self-employment (publicly observed binary output) and hidden firm employment (a flat wage, with individual performance not disclosed to the outside market).222“Hidden” means the market does not see an informative individual-level signal. Inside the firm, managers may observe rich performance signals; we discuss partial public observability in Section 7. The market maintains a public belief about the worker’s talent summarized by a Beta posterior; self-employment updates this public state via observed successes or failures, while hidden employment leaves it unchanged. [^3] Conjugacy keeps the public state finite-dimensional. None of the comparative statics requires Beta priors; any regular one-dimensional prior with monotone likelihood ratio yields the same single-crossing logic. Firms post flat wages. In a benchmark (myopic) case, wages are set equal to the public posterior mean, ignoring the information in application decisions. In the equilibrium case, firms infer from who is expected to apply and post the mean talent of the implied applicant pool.
This setting captures a common choice in modern labor markets: whether to invest in a public track record that travels across jobs, or to accept an opaque wage inside a firm. The trade-off is first-order. Self-employment exposes the worker to risk but creates portable information that improves future outside options; hidden employment insures current consumption but slows market learning. The paper’s contribution is to characterize the dynamic interaction of this insurance–information trade-off with selection into public production and with firms’ inference from applications, delivering sharp comparative statics and testable implications for occupations with portable public histories.
The worker’s problem has a simple structure. Because current public production raises expected utility more for higher-ability workers and because the posted wage is type-independent in a period, the difference between choosing public production and hidden employment is strictly increasing in talent and strictly decreasing in the posted wage. We formalize this single-crossing property in Lemma 1.444The result survives multi-period wages that are independent of contemporaneous individual output and extends to strictly concave $u$ beyond our normalizations. The only knife-edge failure is when current wages do condition on individual output publicly, which restores public learning inside the firm. It implies cutoff policies in the public state and underpins all subsequent results.
Our first main result is an absorbing hidden-employment region. In the benchmark (myopic) case, the posted wage equals the public posterior mean. If a worker below the cutoff takes a hidden job, the public state does not change next period, so the same comparison repeats—hidden employment is absorbing (Proposition 2). The same logic extends to equilibrium posting, where wages account for application-based inference (Proposition 1). Intuition: for sufficiently low public beliefs, the insurance value of a flat wage dominates the informational value of producing another public signal, and because hidden employment generates no public updates, the choice becomes self-confirming.
Our second result establishes existence and uniqueness of the equilibrium cutoff when firms infer from applications. Equilibrium posting ties the wage to acceptance: the posted wage equals the truncated-posterior mean of the types who apply, and the cutoff solves the fixed point. Existence and uniqueness follow from the monotonicity in Lemma 1: the applicant-pool mean is strictly increasing in the cutoff, while indifference is strictly decreasing in the wage (Proposition 4). Intuition: adding marginally stronger applicants raises the equilibrium wage, which makes employment relatively more attractive for slightly weaker types, pinning down a unique fixed point.
Our third result compares equilibrium to the benchmark. Because equilibrium wages price the selected applicant pool, they are (weakly) below benchmark wages at the same public state. Consequently, the equilibrium acceptance cutoff is (weakly) lower (Proposition 5). Intuition: inference from applications corrects the myopic benchmark for adverse selection, compressing wages and tightening self-selection.
The framework yields transparent empirical implications. Early in careers—when a single public outcome moves beliefs more—stronger workers should be more willing to generate public signals; later, absorbing hidden employment becomes more prevalent. Transitions into hidden employment should be more common after runs of low public realizations. When employers price the applicant pool, wage dispersion after observed successes and failures should be compressed relative to a myopic benchmark.
Roadmap. Section 2 reviews related work. Section 3 presents the model. Section 4 characterizes equilibrium under benchmark (myopic) and equilibrium firms and establishes the absorbing-employment result. Section 6 studies comparative statics and testable implications. Section 7 discusses extensions and empirical connections. Section 8 concludes.
2 Related Literature
Our analysis speaks to facts from modern labor markets. Online platforms and open-source ecosystems provide portable, public records of performance; causal evidence shows that public signals substantially improve workers’ subsequent outcomes (Pallais, 2014; Stanton and Thomas, 2016; Horton, 2017). At the same time, much productive activity is not externally visible, and employers gradually learn about workers through noisy, partly private signals (Altonji and Pierret, 2001). By allowing workers to choose between public and hidden production technologies, our framework offers a unified explanation for when we should observe self-employment as a reputational strategy and when we should see long spells of opaque employment with limited outside learning.
We connect classic career-concerns models (Holmström, 1999; Dewatripont et al., 1999a, b; Gibbons and Murphy, 1992) to environments with coexisting public and hidden performance. Relative to that literature, we (i) endogenize the observability choice by letting workers switch between self-employment and hidden firm employment; (ii) characterize an absorbing employment region that rationalizes persistent opacity and slow social learning; and (iii) show how application decisions themselves become signals for equilibrium firms. These elements yield predictions for the incidence and timing of switching, wage dynamics across transparency regimes, and the response to changes in risk exposure or output variance.
Our starting point is the dynamic-career-concerns framework in which market beliefs about talent respond to observed performance (Holmström, 1999). The structure and publicity of information are central to incentives: better information can increase or decrease effort depending on how signals reach the market and how they are aggregated (Dewatripont et al., 1999a, b). We also draw on Gibbons and Murphy (1992), who show that explicit incentive contracts should optimally account for the implicit incentives coming from career concerns. In our setting, the implicit incentives arise predominantly from the worker’s choice of where to produce information (publicly vs. opaquely), rather than from an explicit pay-for-performance scheme.
Our analysis is related to dynamic reputation models in which current actions affect future beliefs (e.g., Board and Meyer-ter-Vehn, 2013). We differ by making the observability of outcomes endogenous through the employment-mode decision. This generates selection on reputation: relatively strong types choose public production to separate, which amplifies the informativeness of public signals; weak types value insurance and hide. The result is an absorbing employment region in which workers optimally stop producing public information—a mechanism that slows social learning even in otherwise transparent markets.
A second thread studies the social value of early-career information and the difficulty firms face in appropriating its benefits. Terviö (2009) shows how markets can under-invest in talent discovery when the returns to information about novice workers are largely external. Empirically, making performance public on online labor markets or via credible third parties causally improves workers’ subsequent outcomes (Pallais, 2014; Stanton and Thomas, 2016; Horton, 2017). Our theory provides conditions under which workers voluntarily seek such transparency by going solo, and conditions under which they rationally prefer opaque employment.
A large empirical literature documents how employers learn about worker productivity over time and the extent to which that learning is public or private (Altonji and Pierret, 2001). We contribute a distinct mechanism: when firm performance is not disclosed externally, the act of applying for such employment becomes informative to equilibrium employers. This echoes broader insights that markets learn not only from realized outcomes but also from selection and participation decisions. In equilibrium, wages and hiring thresholds internalize the informational content of applications even if post-hire outputs remain opaque.
Our applications to open-source contributions and freelancing platforms connect to research on reputation systems and public track records (Bar-Isaac and Tadelis, 2008). In such settings, outside audiences can readily observe some outputs, while many internal tasks remain hidden. The predictions of our model—about switching, absorbing hidden employment for weaker reputations, and wage dynamics across transparency regimes—organize a number of recent empirical facts and suggest new tests, such as how shocks to the visibility of output (e.g., the introduction of new public metrics) shift the incidence of self-employment and the slope of wage–reputation profiles across markets.
Across these strands, our contribution is to endogenize the generation of public information through an employment-mode choice and to analyze how this interacts with firm sophistication in inference. Doing so yields new comparative statics and testable predictions for the distribution of public track records, the clustering of opaque employment spells, and the mapping from reputation to wages when some production is deliberately hidden.
3 Model
Environment
Time is discrete, $t=0,1,2,\ldots$, and the common discount factor is $\delta\in(0,1)$. A single worker has a fixed talent $\theta\in[0,1]$ that is known to the worker and unknown to the outside market. [^5] Allowing for worker learning about own talent is discussed in an extension; here talent is fixed and privately known. In each period the worker chooses between (i) self-employment ($S$), which yields a binary outcome $y_{t}\in\{0,1\}$ with $\Pr(y_{t}=1\mid\theta)=\theta$ that is publicly observed by the market, and (ii) employment at a firm ($E$), which pays a flat wage $w_{t}$ but keeps the worker’s individual outcome hidden from the outside market. In our baseline, the firm does not disclose individual outcomes externally. [^6] In Section 7 we allow the firm to observe a private signal and discuss renegotiation.
Preferences and payoffs
Consumption in period $t$ equals $c_{t}=y_{t}$ under $S$ and $c_{t}=w_{t}$ under $E$ (we normalize the monetary payoff of a success to $1$). The worker has Bernoulli utility $u:\mathbb{R}_{+}\to\mathbb{R}$ that is increasing and concave, with normalization $u(0)=0$ and u(1)=1u(1)=1777This pins units so the expected period utility from self-employment equals the success probability $\theta$. Any positive affine transformation of $u$ would merely re-scale wages; all choices are invariant.. Hence the current utility from $S$ is $\theta$ and from $E$ is $u(w_{t})$. Risk aversion (concavity of $u$) matters through comparison with the certain wage under $E$ and through the value of reputation for future wages.
Beliefs and the public state
[…]
$\mu(\alpha,\beta)\;=\;\frac{\alpha}{\alpha+\beta}.$
A period of hidden employment leaves $(\alpha,\beta)$ unchanged. We write $\mu_{t}\equiv\mu(\alpha_{t},\beta_{t})$.
Firms and wage posting
Each period a competitive fringe of short-lived firms posts a flat wage for applicants. Firms cannot condition the wage on contemporaneous individual performance and do not disclose worker-specific outcomes to the market. [^9] Think unverifiable or multi-task output, team production, or high disclosure costs. Allowing the firm to privately observe a signal and renegotiate does not restore public learning; the absorbing logic only vanishes when the market receives a sufficiently informative public signal from firm employment (Section 7). A firm that hires a worker of talent $\theta$ produces expected revenue $\theta$ (equal to the success probability) and pays wage $w_{t}$. Competition and free entry imply zero expected profits, so the posted wage equals the expected talent of the applicant pool relevant to that firm in that period.
We study two information benchmarks:
  • Naïve firms. Firms ignore the informational content of applying for an opaque firm job. They price off the public state only, posting
    $w^{N}(\alpha,\beta)\;=\;\mu(\alpha,\beta).$ (1)
  • Equilibrium firms. Firms correctly infer that application/acceptance is endogenous and update from the application decision. Let $\Theta^{E}(\alpha,\beta)\subseteq[0,1]$ denote the set of types who apply at state $(\alpha,\beta)$ under the worker’s strategy. Then posted wages satisfy
    $w^{S}(\alpha,\beta)\;=\;\mathbb{E}\big{[}\theta\mid\theta\in\Theta^{E}(\alpha,\beta);\;\alpha,\beta\big{]},$ (2)
[…]
In equilibrium under mild regularity (proved later), the worker’s strategy is monotone in $\theta$ and $\Theta^{E}(\alpha,\beta)$ is an interval $0,\hat{\theta}(\alpha,\beta)]$; then $w^{S}$ is the truncated-posterior mean.
Timing within a period
At the start of period $t$, the public state $(\alpha_{t},\beta_{t})$ is common knowledge. Firms simultaneously post wages as in (1) or (2) given their inference rule. The worker observes $(\alpha_{t},\beta_{t})$ and his own $\theta$ and chooses $a_{t}\in\{S,E\}$. If $a_{t}=S$, the realization $y_{t}\in\{0,1\}$ is drawn and publicly observed, and the state updates to $(\alpha_{t+1},\beta_{t+1})=(\alpha_{t}+1,\beta_{t})$ if $y_{t}=1$ and $(\alpha_{t},\beta_{t}+1)$ if $y_{t}=0$. If $a_{t}=E$, the worker is paid $w_{t}$ and the public state remains $(\alpha_{t+1},\beta_{t+1})=(\alpha_{t},\beta_{t})$.
The worker’s problem
Let $V(\theta;\alpha,\beta)$ denote the worker’s value function. For a given wage policy $w(\alpha,\beta)$ (naïve or equilibrium), the Bellman equation is
$V(\theta;\alpha,\beta)$ $=\max\left\{\,U_{S}(\theta;\alpha,\beta),\;U_{E}(\theta;\alpha,\beta)\,\right\},$ (3)
$U_{S}(\theta;\alpha,\beta)$ $=\theta+\delta\Big{[}\,\theta\,V(\theta;\alpha+1,\beta)+(1-\theta)\,V(\theta;\alpha,\beta+1)\,\Big{]},$ (4)
$U_{E}(\theta;\alpha,\beta)$ $=u\!\left(w(\alpha,\beta)\right)+\delta\,V(\theta;\alpha,\beta).$ (5)
The term $\theta$ in (4) is current utility $\mathbb{E}[u(y_{t})\mid S,\theta]$ under our normalization $u(1)=1$, $u(0)=0$.
Single crossing and cutoff strategies
Define $\Delta(\theta;\alpha,\beta)\equiv U_{S}(\theta;\alpha,\beta)-U_{E}(\theta;\alpha,\beta)$. Under (3)–(5), $\Delta(\theta;\alpha,\beta)$ is strictly increasing in $\theta$: higher $\theta$ raises current expected utility under $S$ and increases the probability of transitioning to the higher-mean state $(\alpha+1,\beta)$. [^11] Formally, $U_{S}$ is strictly increasing in $\theta$, while $U_{E}$ depends on $\theta$ only through the continuation value $V(\theta;\alpha,\beta)$, which is increasing but with slope bounded by $\delta$; the current term in $U_{S}$ delivers the strict single crossing. A complete proof is in Appendix A. Hence there exists a unique cutoff $\hat{\theta}(\alpha,\beta)\in[0,1]$ such that types $\theta\leq\hat{\theta}(\alpha,\beta)$ prefer employment and types $\theta>\hat{\theta}(\alpha,\beta)$ prefer self-employment:
$\theta\leq\hat{\theta}(\alpha,\beta)\;\Longleftrightarrow\;U_{E}(\theta;\alpha,\beta)\geq U_{S}(\theta;\alpha,\beta).$ (6)
Monotonicity also implies $\Theta^{E}(\alpha,\beta)=[0,\hat{\theta}(\alpha,\beta)]$.
Equilibrium
[…]
  1. 1.
    Worker optimality. Given $w(\cdot)$ and beliefs, the worker’s policy is (6) and $V$ solves (3)–(5).
  2. 2.
    Firm optimality and zero profits. In the naïve benchmark, $w=w^{N}$ in (1). In the equilibrium case, wages satisfy $w=w^{S}$ in (2) given the implied applicant set $0,\hat{\theta}(\alpha,\beta)]$.
  3. 3.
    Belief consistency. Beliefs update by Bayes’ rule after public outcomes of $S$ (Beta posterior updates) and remain at $(\alpha,\beta)$ after $E$ (no public signal).
In the equilibrium case, conditions (1) and (2) jointly pin down the fixed point $(\hat{\theta}(\alpha,\beta),w^{S}(\alpha,\beta))$.
Absorbing employment region
A state–type pair $(\alpha,\beta,\theta)$ is said to be absorbing in employment if, conditional on choosing $E$ at $(\alpha,\beta)$, the worker strictly prefers $E$ again at $(\alpha,\beta)$ next period. Because $E$ generates no public signal, the public state remains $(\alpha,\beta)$, so the cutoff $\hat{\theta}(\alpha,\beta)$ is unchanged. The single-crossing property immediately yields:
Proposition 1.
Fix a public state $(\alpha,\beta)$. For any $\theta\leq\hat{\theta}(\alpha,\beta)$, if the worker chooses $E$ at $(\alpha,\beta)$, then $(\alpha,\beta,\theta)$ is absorbing in employment: the worker chooses $E$ in all subsequent periods. For any $\theta>\hat{\theta}(\alpha,\beta)$, the worker rejects $E$ at $(\alpha,\beta)$ and chooses $S$.
Intuitively, for low-reputation public states $(\alpha,\beta)$, the combination of (i) a low posted wage (especially when firms are equilibrium and price the selected applicant pool) and (ii) risk aversion makes the certain wage comparatively attractive for lower types. Because hidden employment produces no new public information, the belief state does not improve, and the same comparison repeats indefinitely. [^12] Formally, the public state $(\alpha,\beta)$ is a controlled Markov state that only moves under self-employment; choosing hidden employment is a self-confirming policy at any state where it is optimal.
Benchmark versus equilibrium pricing
In the naïve benchmark, wages equal the public mean $\mu(\alpha,\beta)$ regardless of application. In the equilibrium case, the wage equals the mean of the truncated posterior on $0,\hat{\theta}(\alpha,\beta)]$. The two benchmarks therefore generate different cutoffs: the equilibrium wage $w^{S}$ incorporates the adverse selection induced by (6), which generally lowers $w$ relative to $w^{N}$ and feeds back into the worker’s choice. Section 4 characterizes the fixed point and shows how pricing sophistication shifts $\hat{\theta}(\alpha,\beta)$ and the size of the absorbing region.
Remarks
(i) The Beta prior delivers a two-dimensional sufficient statistic $(\alpha,\beta)$ and clean updating; none of our qualitative results hinge on conjugacy. (ii) Allowing the current employer to observe a private performance signal and renegotiate wages (without public disclosure) does not restore public learning; Section 7 shows that the key absorbing logic persists. (iii) Off-path beliefs about quitting from firm employment can be parameterized by a single pessimism parameter $\kappa\in[0,1]$ without affecting the baseline results; we return to this in Section 4.
4 Equilibrium and characterization
This section characterizes equilibrium policies and wages under (i) a benchmark with naïve pricing and (ii) equilibrium pricing where firms infer from applications. We also establish the existence of an absorbing employment region and provide comparative statics in patience and risk aversion.
4.1 Benchmark (myopic) wage posting
Setup
In the benchmark (myopic) case, the posted wage depends only on the public state: [^13] Competition and free entry deliver zero expected profits. With a positive profit wedge $\pi$, the analysis goes through with $w^{N}=\mu-\pi$, shifting the cutoff monotonically.
$w^{N}(\alpha,\beta)=\mu(\alpha,\beta)\equiv\frac{\alpha}{\alpha+\beta}.$ (7)
If the worker chooses employment, the public state does not change; if she chooses self-employment, the state updates to $(\alpha+1,\beta)$ after a success and to $(\alpha,\beta+1)$ after a failure. Define
$\begin{split}\mu^{+}(\alpha,\beta)\equiv\mu(\alpha+1,\beta)&=\mu+\frac{\beta}{(\alpha+\beta)(\alpha+\beta+1)},\\ \mu^{-}(\alpha,\beta)\equiv\mu(\alpha,\beta+1)&=\mu-\frac{\alpha}{(\alpha+\beta)(\alpha+\beta+1)}.\end{split}$
Single crossing and cutoff
Let $\Delta(\theta;\alpha,\beta)\equiv U_{S}(\theta;\alpha,\beta)-U_{E}(\theta;\alpha,\beta)$ with $U_{S},U_{E}$ as in (4)–(5). The function $\Delta(\theta;\alpha,\beta)$ is strictly increasing in $\theta$, hence there exists a unique cutoff $\hat{\theta}^{N}(\alpha,\beta)\in[0,1]$ such that types $\theta\leq\hat{\theta}^{N}(\alpha,\beta)$ choose employment and types $\theta>\hat{\theta}^{N}(\alpha,\beta)$ choose self-employment:
$\theta\leq\hat{\theta}^{N}(\alpha,\beta)\quad\Longleftrightarrow\quad U_{E}(\theta;\alpha,\beta)\geq U_{S}(\theta;\alpha,\beta).$ (8)
Proof sketch. $U_{S}$ is strictly increasing in $\theta$ because both the current term and the transition to the more favorable state $(\alpha+1,\beta)$ scale with $\theta$. $U_{E}$ depends on $\theta$ only through the continuation value at the same state $(\alpha,\beta)$, hence has a weaker slope. Therefore $\Delta$ is strictly increasing and crosses zero once. ∎
Absorbing employment
Because employment leaves the public state unchanged, the cutoff at $(\alpha,\beta)$ is time invariant.
Proposition 2.
Fix $(\alpha,\beta)$. If the worker chooses employment at $(\alpha,\beta)$ and $\theta\leq\hat{\theta}^{N}(\alpha,\beta)$, then she chooses employment forever in all subsequent periods. If $\theta>\hat{\theta}^{N}(\alpha,\beta)$, she rejects employment at $(\alpha,\beta)$ and chooses self-employment.
The absorbing region is therefore
$\mathcal{A}^{N}\equiv\{\,(\alpha,\beta,\theta):\theta\leq\hat{\theta}^{N}(\alpha,\beta)\,\}.$
Comparative statics at the benchmark
The cutoff $\hat{\theta}^{N}(\alpha,\beta)$ depends on (i) the posted wage $\mu(\alpha,\beta)$, (ii) the discount factor $\delta$, and (iii) preferences $u(\cdot)$.
Proposition 3.
(i) $\hat{\theta}^{N}(\alpha,\beta)$ is weakly increasing in $\mu(\alpha,\beta)$: a higher public mean raises the wage and expands the employment region. (ii) $\hat{\theta}^{N}(\alpha,\beta)$ is weakly decreasing in $\delta$: greater patience raises the value of generating public information and shrinks the employment region. (iii) Holding $w^{N}$ fixed, if $u$ becomes more concave in the sense of second-order stochastic dominance (i.e., the worker becomes more risk averse), then $\hat{\theta}^{N}(\alpha,\beta)$ weakly increases. [^14] Risk aversion affects the comparison through $u(w^{N})$ in $U_{E}$. With our normalization $u(1)=1$, $u(0)=0$, the current-period utility from self-employment equals $\theta$, while the utility from the certain wage is $u(\mu)\leq\mu$. A more concave $u$ lowers $u(\mu)$ and therefore makes employment less attractive at the same wage; at the equilibrium cutoff, this translates into a weak increase of $\hat{\theta}^{N}$.
Proof sketch. (i) follows from a monotone shift of $U_{E}$ in $w$; (ii) follows because the marginal value of a public signal (the option value embedded in $U_{S}$) scales with $\delta$ while $U_{E}$ holds the state fixed; (iii) is immediate from the comparison $u(\mu)$ vs. $\theta$ at the cutoff. Full proofs appear in Appendix A. ∎
Dynamics across states
As $(\alpha+\beta)$ grows, the one-step drift in beliefs from a public outcome, $|\mu^{\pm}-\mu|$, becomes smaller; this reduces the informational option value of self-employment and expands $\mathcal{A}^{N}$. Thus, late in a career with a long public track record, the incentives to keep producing public signals weaken.
4.2 Equilibrium wage posting and inference from applications
Endogenous applicant pools
With equilibrium firms, wages reflect selection into employment. Given a cutoff policy $\hat{\theta}(\alpha,\beta)$, the applicant pool at state $(\alpha,\beta)$ is $\Theta^{E}(\alpha,\beta)=[0,\hat{\theta}(\alpha,\beta)]$, and the competitive wage equals the pool mean talent under the Beta posterior truncated to this interval:
$w^{S}\!\left(\alpha,\beta;\hat{\theta}\right)\;=\;\mathbb{E}\!\left[\theta\mid\theta\in[0,\hat{\theta}];\;\text{Beta}(\alpha,\beta)\right]\;=\;\frac{B(\hat{\theta};\,\alpha+1,\beta)}{B(\hat{\theta};\,\alpha,\beta)},$ (9)
where $B(c;a,b)\equiv\int_{0}^{c}t^{a-1}(1-t)^{b-1}\,dt$ is the (lower) incomplete beta function.
Fixed point
[…]
$\Delta\!\left(\hat{\theta};\alpha,\beta;\,w^{S}(\alpha,\beta;\hat{\theta})\right)=0,$ (10)
where $\Delta$ is strictly increasing in its first argument.
Lemma 1.
[…]
Proposition 4.
For any $(\alpha,\beta)$ and $\delta\in(0,1)$, there exists a unique cutoff $\hat{\theta}^{S}(\alpha,\beta)\in[0,1]$ solving (10). The corresponding equilibrium wage is $w^{S}(\alpha,\beta;\hat{\theta}^{S})$ from (9). The employment region $\mathcal{A}^{S}\equiv\{(\alpha,\beta,\theta):\theta\leq\hat{\theta}^{S}(\alpha,\beta)\}$ is absorbing.
[…]
Naïve versus equilibrium
Equilibrium pricing internalizes adverse selection: $w^{S}(\alpha,\beta;\hat{\theta})\leq w^{N}(\alpha,\beta)$ with strict inequality whenever $\hat{\theta}<1$. Lower wages tighten the employment condition and shift the cutoff down.
Proposition 5.
For any $(\alpha,\beta)$, $\hat{\theta}^{S}(\alpha,\beta)\leq\hat{\theta}^{N}(\alpha,\beta)$, with a strict inequality when $\hat{\theta}^{N}(\alpha,\beta)<1$.
Proof sketch. For a given state, $w^{S}\leq w^{N}$ because the applicant pool mean does not exceed the unconditional posterior mean. Since $U_{E}$ is increasing in the wage and $\Delta$ is increasing in $\theta$, the indifference point shifts left. ∎
4.3 Comparative statics and testable implications
Patience
Increasing $\delta$ raises the option value of generating public signals. In both pricing regimes, $\hat{\theta}(\alpha,\beta)$ is weakly decreasing in $\delta$. [^16] At the equilibrium regime this follows from the envelope property of $V$ and the fact that employment leaves the public state unchanged while self-employment stochastically moves it toward higher $\mu$ with probability $\theta$.
Risk aversion
Holding the wage schedule fixed, greater concavity of $u$ weakly increases $\hat{\theta}$ (employment becomes less attractive at a given wage because $u(w)$ falls). With equilibrium pricing, an additional (general-equilibrium) effect operates through composition: when fewer types apply, the applicant pool is stronger, which increases $w^{S}$ and partially offsets the direct effect. Overall, $\hat{\theta}^{S}$ remains weakly increasing in risk aversion under standard regularity (Appendix C).
Public-track-record length
As $(\alpha+\beta)$ increases, the one-step drift $|\mu^{\pm}-\mu|$ shrinks, reducing the informational benefit of an additional public signal. Hence $\hat{\theta}(\alpha,\beta)$ is weakly increasing in $(\alpha+\beta)$, and the absorbing employment region expands with career length.
Cross-sectional predictions
The model predicts: (i) stronger reputations (higher $\mu$) are associated with higher acceptance of employment offers when pricing is naïve, but with a muted pattern under equilibrium pricing; (ii) transitions from self-employment to opaque employment are more common after runs of low public performance; (iii) wage offers by equilibrium firms are lower, on average, than naïve benchmark wages at the same public state, reflecting adverse selection.
A note on quitting and off-path beliefs
If the public observes quits from opaque employment, off-path beliefs can be parameterized by $\kappa\in[0,1]$ measuring the pessimism attached to an unexplained quit (Section 7). For a range of $\kappa$, the absorbing region characterized above is unchanged, because beliefs do not update while the worker remains in opaque employment.
5 Three-period illustration
To visualize the mechanisms, consider a finite horizon with three decision dates $t=0,1,2$ (the last action occurs at $t=2$). The worker discounts with factor $\delta\in(0,1)$. For transparency and closed forms, suppose the prior at $t=0$ is uniform, $\text{Beta}(1,1)$, and the period utility over certain consumption $x\in[0,1]$ is CRRA on $[0,1]$, $u(x)=x^{\rho}$ with $\rho\in(0,1]$; [^17] At $\rho=1$ (risk neutrality) the period utility comparison collapses to means; the qualitative difference between benchmark and equilibrium posting still comes entirely from application-based inference. recall $u(1)=1,u(0)=0$, so $\mathbb{E}[u(y_{t})\mid S,\theta]=\theta$.
5.1 Last period ($t=2$)
At $t=2$ there is no continuation value. The worker compares $S$ (utility $\theta$) to $E$ (utility $u(w_{2})$), so the cutoff is $\hat{\theta}_{2}=u(w_{2})$.
Naïve pricing
Wages equal the public posterior mean at the current state:
$w^{N}_{2}(\alpha,\beta)=\mu(\alpha,\beta)=\frac{\alpha}{\alpha+\beta}.$
From the uniform prior, the three relevant states and wages are
$(\alpha,\beta)\in\{(1,1),(2,1),(1,2)\},\quad w^{N}_{2}\in\Big{\{}\tfrac{1}{2},\tfrac{2}{3},\tfrac{1}{3}\Big{\}}.$
Hence the last-period cutoffs under naïve pricing are
$\hat{\theta}^{N}_{2}(1,1)=(\tfrac{1}{2})^{\rho},\qquad\hat{\theta}^{N}_{2}(2,1)=(\tfrac{2}{3})^{\rho},\qquad\hat{\theta}^{N}_{2}(1,2)=(\tfrac{1}{3})^{\rho}.$
Equilibrium pricing
Firms infer from applications. With a cutoff policy $\hat{\theta}$, the last-period wage equals the truncated-posterior mean on $0,\hat{\theta}$]. For $\text{Beta}(\alpha,\beta)$ this mean is
$w^{S}_{2}(\alpha,\beta;\hat{\theta})\;=\;\frac{B(\hat{\theta};\,\alpha+1,\beta)}{B(\hat{\theta};\,\alpha,\beta)},$
with $B(\cdot)$ the incomplete beta function. The cutoff solves the fixed point
$\hat{\theta}^{S}_{2}(\alpha,\beta)=u\!\left(w^{S}_{2}(\alpha,\beta;\hat{\theta}^{S}_{2}(\alpha,\beta))\right).$
From the uniform prior:
$(\alpha,\beta)=(1,1):\quad$ $w^{S}_{2}(\hat{\theta})=\tfrac{1}{2}\hat{\theta}\ \Rightarrow\ \hat{\theta}^{S}_{2}(1,1)=\Big{(}\tfrac{\hat{\theta}}{2}\Big{)}^{\rho}\ \Rightarrow\ \hat{\theta}^{S}_{2}(1,1)=2^{-\rho/(1-\rho)}\quad(\rho<1),$
$(\alpha,\beta)=(2,1):\quad$ $w^{S}_{2}(\hat{\theta})=\tfrac{2}{3}\hat{\theta}\ \Rightarrow\ \hat{\theta}^{S}_{2}(2,1)=\Big{(}\tfrac{2}{3}\hat{\theta}\Big{)}^{\rho}\ \Rightarrow\ \hat{\theta}^{S}_{2}(2,1)=\left(\tfrac{2}{3}\right)^{\rho/(1-\rho)},$
$(\alpha,\beta)=(1,2):\quad$ $w^{S}_{2}(\hat{\theta})=\frac{\hat{\theta}-\frac{2}{3}\hat{\theta}^{2}}{\,2-\hat{\theta}\,}\ \Rightarrow\ \hat{\theta}^{S}_{2}(1,2)\text{ solves }\ \hat{\theta}=\Big{(}\frac{\hat{\theta}-\frac{2}{3}\hat{\theta}^{2}}{\,2-\hat{\theta}\,}\Big{)}^{\rho}.$
These solutions are interior for $\rho\in(0,1)$ and satisfy $\hat{\theta}^{S}_{2}(\alpha,\beta)\leq\hat{\theta}^{N}_{2}(\alpha,\beta)$, with equality only if $\hat{\theta}=1$ (no selection).
Numerical illustration
With $\rho=\tfrac{1}{2}$ (square-root) and, say, $\delta=0.95$ for the next subsection, the last-period cutoffs are:
$\hat{\theta}^{N}_{2}(1,1)\approx 0.707,\quad\hat{\theta}^{N}_{2}(2,1)\approx 0.817,\quad\hat{\theta}^{N}_{2}(1,2)\approx 0.577;$
$\hat{\theta}^{S}_{2}(1,1)=2^{-1}=0.5,\quad\hat{\theta}^{S}_{2}(2,1)=(2/3)^{1}=0.667,\quad\hat{\theta}^{S}_{2}(1,2)\approx 0.451.$
Thus sophistication lowers wages and cutoffs (relative to naïve), especially at low-belief states.
5.2 Middle period ($t=1$) and backward induction
Let $V_{2}(\theta;\alpha,\beta)=\max\{\theta,\ u(w_{2}(\alpha,\beta))\}$ denote the last-period value. At $t=1$ the worker compares
$U_{S}^{(1)}(\theta;\alpha,\beta)=\theta+\delta\Big{[}\,\theta\,V_{2}(\theta;\alpha+1,\beta)+(1-\theta)\,V_{2}(\theta;\alpha,\beta+1)\,\Big{]},$
$U_{E}^{(1)}(\theta;\alpha,\beta)=u(w_{1}(\alpha,\beta))+\delta\,V_{2}(\theta;\alpha,\beta),$
with $w_{1}=w^{N}$ or $w^{S}(\cdot;\hat{\theta}_{1})$ depending on the pricing regime. Strict single crossing in $\theta$ implies a unique middle-period cutoff $\hat{\theta}_{1}(\alpha,\beta)$.
Naïve pricing, uniform prior
At state $(1,1)$ we have $w^{N}_{1}=1/2$ and
$\begin{split}V_{2}(\theta;2,1)&=\max\{\theta,(2/3)^{\rho}\},\\ V_{2}(\theta;1,2)&=\max\{\theta,(1/3)^{\rho}\},\\ V_{2}(\theta;1,1)&=\max\{\theta,(1/2)^{\rho}\}.\end{split}$
The cutoff $\hat{\theta}^{N}_{1}(1,1)$ solves
$\hat{\theta}+\delta\Big{[}\hat{\theta}\,V_{2}(\hat{\theta};2,1)+(1-\hat{\theta})\,V_{2}(\hat{\theta};1,2)\Big{]}\;=\;(1/2)^{\rho}+\delta\,V_{2}(\hat{\theta};1,1).$
With $\rho=\tfrac{1}{2}$ and $\delta=0.95$, this yields [^18] In the relevant region $\hat{\theta}\in[0.577,0.707]$, $V_{2}(\hat{\theta};2,1)=(2/3)^{1/2}$, $V_{2}(\hat{\theta};1,2)=\hat{\theta}$, and $V_{2}(\hat{\theta};1,1)=(1/2)^{1/2}$, reducing the equation to a quadratic.
$\hat{\theta}^{N}_{1}(1,1)\approx 0.656.$
Equilibrium pricing, uniform prior
Now $w^{S}_{1}(1,1;\hat{\theta}_{1})=\mathbb{E}[\theta\mid\theta\leq\hat{\theta}_{1}]=\hat{\theta}_{1}/2$, so $u(w^{S}_{1})=\big{(}\hat{\theta}_{1}/2\big{)}^{\rho}$. The relevant last-period values use the state-specific equilibrium cutoffs from above:
$\begin{split}V_{2}^{S}(\theta;2,1)&=\max\{\theta,\ \hat{\theta}^{S}_{2}(2,1)\},\\ V_{2}^{S}(\theta;1,2)&=\max\{\theta,\ \hat{\theta}^{S}_{2}(1,2)\},\\ V_{2}^{S}(\theta;1,1)&=\max\{\theta,\ \hat{\theta}^{S}_{2}(1,1)\}.\end{split}$
At $(1,1)$ with $\rho=\tfrac{1}{2}$ and $\delta=0.95$, the fixed-point indifference equation becomes
$\hat{\theta}+\delta\Big{[}\hat{\theta}\,\max\{\hat{\theta},\,2/3\}+(1-\hat{\theta})\,\max\{\hat{\theta},\,0.451\}\Big{]}=\sqrt{\hat{\theta}/2}\;+\;\delta\,\max\{\hat{\theta},\,0.5\}.$
The unique solution is
$\hat{\theta}^{S}_{1}(1,1)\approx 0.42,$
substantially below $\hat{\theta}^{N}_{1}(1,1)\approx 0.656$, reflecting the adverse-selection discount in equilibrium wages.
Takeaways
Two features stand out. First, absorbing hidden employment: once a worker with $\theta\leq\hat{\theta}_{t}(\alpha,\beta)$ accepts $E$, the public state freezes and the inequality repeats, so employment is absorbing at that state. Second, pricing sophistication compresses wages and tightens self-selection: cutoffs fall ($\hat{\theta}^{S}\leq\hat{\theta}^{N}$), so a smaller set accepts employment early; mid-types instead generate public signals via $S$.
With $u(x)=\sqrt{x}$ and a uniform prior, equilibrium wage posting compresses the success–failure gap at $t=1$: benchmark wages are $2/3$ (after success) vs. $1/3$ (after failure), while equilibrium wages are about $0.407$ vs. $0.177$ under our calibration—a substantial winner’s-curse correction.
Acceptance-set evolution (equilibrium posting).
Under equilibrium wage posting, the applicant pool each date is $\Theta^{E}_{t}(\alpha,\beta)=[0,\hat{\theta}_{t}(\alpha,\beta)]$. Starting from $(1,1)$ at $t=1$, a public success (failure) moves the state to $(2,1)$ (resp. $(1,2)$), which induces distinct acceptance sets $0,\hat{\theta}_{1}(2,1)]$ and $0,\hat{\theta}_{1}(1,2)]$ and thus distinct wages $w^{E}_{1}(2,1)=\mathbb{E}[\theta\mid\theta\leq\hat{\theta}_{1}(2,1)]$ and $w^{E}_{1}(1,2)=\mathbb{E}[\theta\mid\theta\leq\hat{\theta}_{1}(1,2)]$. The same logic applies at $t=2$ for $(\alpha,\beta)\in\{(1,1),(2,1),(1,2)\}$ with $\hat{\theta}_{2}(\alpha,\beta)$ solving the corresponding fixed point. This makes explicit how acceptance sets and wages evolve along each branch of the public history.
6 Empirical implications
Patience
Increasing $\delta$ expands the option value of public signals. In both pricing regimes, the cutoff $\hat{\theta}(\alpha,\beta)$ is weakly decreasing in $\delta$: more patient workers are more willing to forgo insurance to refresh reputation.
Risk aversion
Holding the wage schedule fixed, greater concavity of $u$ (on $[0,1]$ with $u(0)=0,u(1)=1$) raises the utility of a given certain wage relative to a risky Bernoulli payoff with the same mean. Hence $\hat{\theta}(\alpha,\beta)$ is weakly increasing in risk aversion: more types accept employment at a given public state. Under equilibrium pricing, a composition effect operates (fewer applicants strengthen the pool and raise $w^{S}$), which reinforces the shift.
Public-track-record length
As $(\alpha+\beta)$ grows, a single public outcome moves $\mu$ by less ($|\mu^{\pm}-\mu|$ shrinks), reducing the informational benefit of $S$. Thus $\hat{\theta}(\alpha,\beta)$ is weakly increasing in $(\alpha+\beta)$; absorbing hidden employment is more prevalent late in careers.
Transparency premia across occupations
Occupations with portable public histories (open-source, contests, platforms) should exhibit (i) higher early self-employment/portfolio-building by stronger types, (ii) sharper wage gains after publicly observed successes, and (iii) more frequent switches back to opaque employment following poor public realizations.
Application-based inference
With equilibrium pricing, application choices are informative. Relative to naïve environments (e.g., where HR ignores selection), we predict (i) lower average wages for otherwise identical public states, (ii) fewer mid-type applicants, and (iii) stronger sorting on outside reputations.
Empirical probes
The model yields simple tests: (i) event studies of wage paths after public successes/failures for workers who remain outside versus those who enter opaque employment; (ii) cross-occupation comparisons of the share of opaque spells by public-reputation quantiles; (iii) within-platform evidence that mid-quantile workers are more likely to keep generating public signals when employers are more “equilibrium” (e.g., settings with better screening or more experienced recruiters).
7 Discussion and extensions
This section discusses how our mechanisms interact with (i) worker learning about own type, (ii) partial public observability inside firms, and (iii) simple screening menus. The baseline insights—a cutoff policy with an absorbing hidden-employment region and selection into public production—are robust.
7.1 Worker learning about own type
In the baseline, $\theta$ is fixed and privately known. Here the worker starts with a private prior $H_{0}$ and updates after each period using privately observed outcomes; public outcomes of $S$ update both $H_{t}$ and the market’s $F_{t}$, while $E$ updates only $H_{t}$. Let $m_{t}\equiv\mathbb{E}[\theta\mid\text{worker's private history}]$ and $\mu_{t}\equiv\mathbb{E}[\theta\mid\text{public history}]$.
Remark 1.
Fix a public state $(\alpha,\beta)$ (summarized by $\mu_{t}$). For any wage schedule and $\delta\in(0,1)$, there exists a unique cutoff $\hat{m}(\alpha,\beta)$ such that the worker chooses $E$ iff $m_{t}\leq\hat{m}(\alpha,\beta)$; the map $m\mapsto U_{S}(m;\alpha,\beta)-U_{E}(m;\alpha,\beta)$ is strictly increasing.
The absorbing logic extends: if $E$ is optimal at $(\mu_{t},m_{t})$, then the public state remains $(\alpha,\beta)$ and the same comparison repeats. [^19] Private updating under $E$ can move $m_{t+1}$, but with flat wages independent of private realizations, there is a neighborhood of private beliefs for which absorption persists. Public learning is restored only if the market sees an informative signal from $E$. Private learning under $E$ may move $m_{t+1}$ but does not overturn absorption locally with flat wages. Under equilibrium posting (application inference), the cutoff $\hat{m}(\alpha,\beta)$ shifts down due to the adverse-selection discount.
7.2 Partial public observability inside firms
Suppose that under $E$ the market observes a coarse signal $z_{t}\in\{0,1\}$ with informativeness $\phi\in[0,1)$:
$\Pr(z_{t}=1\mid\theta,E)=\phi\,\theta+(1-\phi)\,\bar{z}.$
For sufficiently small $\phi$, the equilibrium still features an absorbing hidden-employment region; as $\phi$ increases, the region shrinks and vanishes only when $E$ generates near-parity information with $S$.
7.3 Simple screening menus
Allow firms to post menus $(w,p)$ where $p$ is the intensity of public-facing tasks (or disclosure). Under competitive posting and zero profits, menus cannot separate types beyond the entry cutoff: raising $p$ to attract higher types requires lowering $w$, which erodes the insurance value and pushes those types back to $S$. The main margin remains entry vs. self-employment; menus tilt the cutoff but do not overturn absorption.
8 Conclusion
This paper studies career concerns when some production is publicly observable and some is opaque. A risk-averse worker chooses each period between self-employment, which generates public signals about talent, and hidden firm employment, which pays a flat wage and produces no public information. Firms post wages either by benchmarking myopically on the public state or, in equilibrium, by inferring from applications and pricing the selected applicant pool. The worker’s problem exhibits single crossing, so behavior is governed by a cutoff policy in the public state.
Three forces organize the results. First, an insurance–information trade-off: hidden employment offers insurance but suppresses public learning and future outside options. Second, selection: relatively stronger types opt to produce public signals, making observed outcomes more informative. Third, inference from applications: equilibrium wage posting discounts for adverse selection in the applicant pool. Together these forces generate an absorbing hidden-employment region—a set of public beliefs for which workers optimally accept hidden employment forever—alongside switching back to public production when reputation is sufficiently strong.
We deliver a simple characterization under both posting regimes. In the benchmark (myopic) case, wages equal the public posterior mean and the absorbing region expands with risk aversion and with career length (as single outcomes move beliefs less). Under equilibrium posting, a fixed point links the acceptance cutoff to the truncated-posterior wage; existence and uniqueness follow from monotonicity of the truncated mean and single crossing in the worker’s problem. Pricing sophistication compresses wages relative to the benchmark at a given public state and tightens self-selection, shifting the cutoff down.
The framework yields transparent empirical implications. Early in careers, when public signals move beliefs more, workers are more willing to generate public information; later, absorbing hidden employment becomes more prevalent. After public successes (failures), the benchmark predicts wide wage dispersion, whereas equilibrium posting corrects for winner’s-curse effects through application-based inference, compressing the success–failure gap. Across occupations, environments with portable public histories (e.g., open portfolios, platform ratings, contest leaderboards) should feature stronger selection into public production by higher-ability workers and sharper wage responses to public outcomes. [^20] Examples include code and data science (open repositories), academic research (working papers and preprints), design (public portfolios), and online freelancing (platform ratings).
We showed that the main insights are robust. Allowing the worker to learn about own type preserves the cutoff structure (in private beliefs) and the absorbing logic because hidden employment freezes the public state. Introducing partial public observability inside firms shrinks, but does not eliminate, the absorbing region unless internal signals become nearly as informative as public production. Letting firms offer simple $(w,p)$ menus shifts the cutoff but does not overturn the central margin: the entry decision mediates the trade-off between insurance and reputation.
Two design lessons follow. First, organizational transparency is a lever: policies that selectively externalize performance (client-visible tasks, standardized public metrics) can substitute for self-employment as a reputational technology, [^21] For instance, making some client KPIs or review summaries public, adopting external code review, or giving structured, verifiable titles that map to outside-market signals. reducing the absorbing hidden-employment region and speeding market learning. Second, marketplace curation (platform labels, verified histories) changes the relative returns to public signals and therefore the equilibrium mix of public versus hidden production.
Naturally, the model is stylized. We abstract from multi-tasking, multi-dimensional ability, strategic project choice, and richer long-term contracts, and we treat firms as short-lived competitors rather than repeat principals. Extending the analysis to relational contracts, internal renegotiation with private performance signals, or general priors beyond the Beta family would sharpen the link to personnel and organizational economics. On the empirical side, the predictions invite structural or event-study tests on settings with observable public tracks and opaque firm spells (e.g., platform markets, open-source ecosystems, or occupations with client-visible KPIs).
By tying a classic career-concerns logic to environments with coexisting public and hidden performance, the paper provides a compact theory of when employment becomes absorbing, how application-based inference reshapes wages, and why the returns to public production are largest early in careers. We hope this framework proves useful for organizing empirical work on modern reputation systems and for guiding organizational choices about transparency and the design of early-career opportunities.
Acknowledgements
We thank Emiliano Catonini and Steven Kivinen for very helpful comments and encouragement on earlier drafts and related work. We are grateful to seminar and workshop participants at our home institutions and elsewhere for suggestions that improved the paper. All remaining errors are our own.
Appendix A: Proofs
Proof of Lemma 1.
(1) With
$U_{S}(\theta;\alpha,\beta)=\theta+\delta\{\theta V(\theta;\alpha{+}1,\beta)+(1{-}\theta)V(\theta;\alpha,\beta{+}1)\}$
[…]
Proof of Proposition 1.
If $\theta\leq\hat{\theta}(\alpha,\beta)$ then $\Delta(\theta;\alpha,\beta,w(\alpha,\beta))\leq 0$. Employment leaves $(\alpha,\beta)$ (and hence $w(\alpha,\beta)$) unchanged, so the same inequality holds next period. By induction, employment is repeatedly optimal. ∎
Proof of Proposition 4.
[…]
Proof of Proposition 5.
At any $(\alpha,\beta)$, $w^{S}(\alpha,\beta;\hat{\theta})\leq w^{N}(\alpha,\beta)$ (strict when $\hat{\theta}<1$) because truncation to $0,\hat{\theta}$] lowers the posterior mean. Since $\Delta$ is strictly decreasing in $w$, the unique zero under $w^{S}$ lies weakly to the left of the zero under $w^{N}$. ∎
Appendix B: Details for truncated Beta and the fixed point
Truncated Beta derivatives.
With $f(t)\propto t^{\alpha-1}(1-t)^{\beta-1}$, $F(c)=\int_{0}^{c}f$ and $A(c)=\int_{0}^{c}tf$, we have $m(c)=A/F$ and
$m^{\prime}(c)=\frac{f(c)}{F(c)}\,[c-m(c)]>0,\qquad\frac{d}{dc}\big{(}u(m(c))\big{)}=u^{\prime}(m(c))\,m^{\prime}(c)>0.$
These identities imply the monotonicity used in the existence/uniqueness proof and are convenient for comparative statics.
Equilibrium mapping.
At any $(\alpha,\beta)$, define $\phi(w)$ by $\Delta(\phi(w);\alpha,\beta,w)=0$. By the implicit function theorem,
$\phi^{\prime}(w)=-\frac{\partial\Delta/\partial w}{\partial\Delta/\partial\theta}<0,$
[…]
Appendix C: Comparative statics with worker learning and partial observability
Worker learning
Let the worker hold a private posterior $H_{t}$ with mean $m_{t}$. Replace $\theta$ in (4) with $m_{t}$ in the current term and in the transition probabilities. The mapping $m\mapsto U_{S}(m)-U_{E}(m)$ is strictly increasing by the same argument as in Lemma 1, so a unique cutoff $\hat{m}(\alpha,\beta)$ exists. Since employment freezes the public state $(\alpha,\beta)$, the same comparison repeats next period; local absorption follows.
Partial public observability
If $E$ generates a Bernoulli public signal with informativeness $\phi\in[0,1)$, posterior updating at $E$ weakly increases the option value of employment. For sufficiently small $\phi$ (bounded away from 1), the signal is too noisy to overturn absorption: the wage still prices the selected pool and the single-crossing comparison is preserved. As $\phi\uparrow 1$, the absorbing region shrinks and vanishes in the limit where $E$ is as informative as $S$.