### v3

# Self-Employment as a Signal:

Abstract
We study a dynamic labor market in which a risk-averse worker with career concerns chooses each period between self-employment, which generates publicly observed binary output, and employment at a firm, which pays a flat wage but keeps individual performance hidden from the outside market. The worker values future opportunities through a reputation for talent, understood as the public belief used by firms to set wages. Firms may be myopic, pricing only off public beliefs, or sophisticated, inferring from who is expected to apply. Three forces organize the results: an insurance–information trade-off, selection by talent, and inference from application decisions. A self-confirming absorbing hidden-employment region arises naturally when the value of insurance dominates the value of additional public signals, so some workers optimally stop generating public information. At any public belief there is a unique talent cutoff governing selection into self-employment, and sophisticated firms, by internalizing application-based inference, compress wages and induce different patterns of public information production than myopic firms. The framework yields testable predictions for the prevalence of self-employment, switching into opaque jobs, and wage dynamics across markets with different degrees of performance transparency.
keywords:
Career concerns , reputation , signaling , self-employment , transparency , labor-market learning
JEL:
J24 , D82 , L26 , D83 , J31
journal: Working Paper
1 Introduction
Markets often see some but not all of what workers do. A developer’s open-source commits, a freelancer’s platform ratings, or a salesperson’s client-visible metrics are public; most day-to-day contributions inside firms are not. We study a dynamic environment in which a risk-averse worker cares about both current consumption and a reputation for talent that shapes future wages. Each period the worker chooses between self-employment (publicly observed binary output) and hidden222“Hidden” means the market does not see an informative individual‐level signal. Inside the firm, managers may observe rich performance signals; we discuss partial public observability in Section 6. firm employment (a flat wage, with individual performance not disclosed to the outside market). The market maintains a public belief about the worker’s talent summarized by a Beta posterior; self-employment updates this public state via observed successes or failures, while hidden employment leaves it unchanged. [^3] Conjugacy keeps the public state finite‐dimensional. None of the comparative statics requires Beta priors; any regular one‐dimensional prior with monotone likelihood ratio yields the same single‐crossing logic. Firms post flat wages. In a benchmark case, wages are set equal to the public posterior mean, ignoring the information in application decisions. When firms are sophisticated, they infer from who is expected to apply and post the mean talent of the implied applicant pool.
We embed this choice in a stationary, infinite-horizon environment with overlapping generations of workers. Time is discrete and there is a constant probability that any given worker exits the market in each period; exiting workers are replaced by new entrants drawn from the same prior. Firms and outside employers do not observe calendar age or total tenure, only the worker’s public record of self-employment outcomes. In equilibrium, wages and cutoff strategies are therefore time-invariant functions of the public state, and our comparative statics can be interpreted as describing steady-state differences across labor markets. [^4] This infinite-horizon, overlapping-generations formulation follows the spirit of Pei, (2025)’s analysis of stationary reputational environments.
This setting captures a common choice in modern labor markets: whether to invest in a public track record that travels across jobs, or to accept an opaque wage inside a firm. Equivalently, one can read the model as a worker deciding how much effort to allocate to public projects—open-source contributions, contests, portfolio pieces, client-visible work—versus hidden projects inside the firm whose outcomes are not externally observable. The trade-off is first-order. Self-employment exposes the worker to risk but creates portable information that improves future outside options; hidden employment insures current consumption but slows market learning. The paper’s contribution is to show how selection into self-employment versus hidden jobs jointly shapes both wages and the amount of public information produced about talent, under naïve and sophisticated wage-setting regimes, and to characterize the resulting comparative statics and testable implications for occupations with portable public histories.
The worker’s problem has a simple structure. Because current public production raises expected utility more for higher-ability workers and because the posted wage is type-independent in a period, the difference between choosing public production and hidden employment is strictly increasing in talent and strictly decreasing in the posted wage. We formalize this single-crossing property in Lemma 1.555The result survives multi‐period wages that are independent of contemporaneous individual output and extends to strictly concave $u$ beyond our normalizations. The only knife‐edge failure is when current wages do condition on individual output publicly, which restores public learning inside the firm. It implies cutoff policies in the public state and underpins all subsequent results.
First, we characterize a self-confirming absorbing hidden-employment region in the state space. For sufficiently low public beliefs about the worker’s talent, the flat wage offered by firms dominates the option value of generating further public signals, so the worker optimally chooses hidden employment; because that choice produces no public updates, the worker then remains in this region going forward. The size and shape of this region respond in intuitive ways to primitives such as risk aversion, patience, and the informativeness of public performance: it expands when workers are more risk averse or opaque wages are generous, and shrinks when workers are more patient, better insured, or operate in more transparent labor markets.
A second key implication is the existence and uniqueness of a well-behaved cutoff: at each public state there is a single talent threshold separating workers who self-employ (to generate or maintain public signals) from those who choose the firm’s insurance, and standard monotonicity arguments ensure that this fixed point is uniquely pinned down. Existence and uniqueness follow from the monotonicity in Lemma 1: the applicant-pool mean is strictly increasing in the cutoff, while indifference is strictly decreasing in the wage (Proposition 4). Intuition: adding marginally stronger applicants raises the equilibrium wage, which makes employment relatively more attractive for slightly weaker types, pinning down a unique fixed point.
We also compare sophisticated pricing to the myopic benchmark: because sophisticated firms internalize adverse selection from application decisions, they offer lower wages than naïve firms, compress the payoff to additional public signals, and tighten self-selection into self-employment versus hidden jobs. Methodologically, our stationary infinite-horizon formulation is in the spirit of Pei, (2025), who studies reputational dynamics with endogenous records in a doubly infinite horizon, and conceptually our hidden-effort extension connects the worker’s observability choice to the work–shirk dynamics in dynamic reputation models such as Board and Meyer-ter Vehn, (2013).
The framework yields transparent empirical implications. Early in careers—when a single public outcome moves beliefs more—stronger workers should be more willing to generate public signals; later, absorbing hidden employment becomes more prevalent. Transitions into hidden employment should be more common after runs of low public realizations. When employers price the applicant pool, wage dispersion after observed successes and failures should be compressed relative to a myopic benchmark.
Roadmap. Section 2 reviews related work. Section 3 presents the model. Section 4 characterizes equilibrium under myopic and sophisticated wage setting and establishes the absorbing-employment result. Section 5 develops comparative statics and empirical implications. Section 6 discusses extensions, policy implications, and alternative interpretations. Section 7 concludes.
2 Related Literature
Our analysis speaks to facts from modern labor markets. Online platforms and open-source ecosystems provide portable, public records of performance; causal evidence shows that public signals substantially improve workers’ subsequent outcomes (Pallais,, 2014; Stanton and Thomas,, 2016; Horton,, 2017). At the same time, much productive activity is not externally visible, and employers gradually learn about workers through noisy, partly private signals (Altonji and Pierret,, 2001). By allowing workers to choose between public and hidden production technologies, our framework offers a unified explanation for when we should observe self-employment as a reputational strategy and when we should see long spells of opaque employment with limited outside learning.
We connect classic career-concerns models (Holmström,, 1999; Dewatripont et al., 1999a, ; Dewatripont et al., 1999b, ; Gibbons and Murphy,, 1992) to environments with coexisting public and hidden performance. Relative to that literature, we (i) endogenize the observability choice by letting workers switch between self-employment and hidden firm employment; (ii) characterize an absorbing employment region that rationalizes persistent opacity and slow social learning; and (iii) show how application decisions themselves become signals for sophisticated firms. These elements yield predictions for the incidence and timing of switching, wage dynamics across transparency regimes, and the response to changes in risk exposure or output variance.
Our starting point is the dynamic-career-concerns framework in which market beliefs about talent respond to observed performance (Holmström,, 1999). The structure and publicity of information are central to incentives: better information can increase or decrease effort depending on how signals reach the market and how they are aggregated (Dewatripont et al., 1999a, ; Dewatripont et al., 1999b, ). We also draw on Gibbons and Murphy, (1992), who show that explicit incentive contracts should optimally account for the implicit incentives coming from career concerns. In our setting, the implicit incentives arise predominantly from the worker’s choice of where to produce information (publicly vs. opaquely), rather than from an explicit pay-for-performance scheme.
Our analysis is related to dynamic reputation models in which current actions affect future beliefs (e.g., Board and Meyer-ter Vehn,, 2013). We differ by making the observability of outcomes endogenous through the employment-mode decision. This generates selection on reputation: relatively strong types choose public production to separate, which amplifies the informativeness of public signals; weak types value insurance and hide. The result is an absorbing employment region in which workers optimally stop producing public information—a mechanism that slows social learning even in otherwise transparent markets.
More broadly, our model is close in spirit to recent work on dynamic career concerns with rich information environments. (Pei,, 2025) analyzes community enforcement with endogenous records in anonymous matching: players can add or erase signals from their histories, and the resulting monotonicity constraints on continuation values sharply restrict cooperation when lifetimes are very long. Our worker’s choice between generating new public signals and freezing the record through hidden employment is a labor-market analogue of this record-design margin. Other recent dynamic signaling models with career concerns emphasize experimentation and complex information flows (e.g. Thomas,, 2019; Kaniel and Orlov,, 2025). On the empirical side, Abou El-Komboz and Goldbeck, (2025) show that developers’ career concerns causally increase open-source contributions, while Belenzon and Schankerman, (2015) document sorting and signaling motives in open innovation. These findings motivate our focus on self-employment and public projects as a reputational technology and on how market transparency shapes the supply of public information about talent.
A large literature studies dynamic careers in internal labor markets. In Gibbons and Waldman, (1999), firms learn about worker ability from internal output and adjust job assignments and wages over time; Farber and Gibbons, (1996) develop a dynamic model of learning and wage dynamics in a competitive labor market; and Gibbons and Waldman, (2006); Waldman, (2012) survey theory and evidence on careers in organizations. In these models, performance is effectively observed by the relevant decision maker (the firm), so learning is primarily about who to promote or retain. By contrast, we focus on environments where some production is deliberately hidden from the outside market, so that the worker’s choice between public and hidden work jointly determines both the information set of future employers and the mapping from reputation into wages. Our absorbing hidden-employment region can be read as a simple mechanism through which careers become opaque and social learning slows down, even when performance is transparent inside firms.
Related empirical work studies learning and selection in online labor markets and digital platforms. Leung, (2018) shows how employers update beliefs about worker quality from their own past hiring experiences in a large online freelance market, while Kokkodis and Ipeirotis, (2014) document how platform ratings and skill signals shape demand and match quality. These papers emphasize that observed performance and feedback in digital environments are central inputs into hiring and pricing decisions. Our framework complements this evidence by modeling a worker’s choice to invest in public versus hidden performance as a reputational technology, and by characterizing how different wage-setting regimes (naïve versus sophisticated) affect both selection into public work and the amount of public information generated about talent.
A second thread studies the social value of early-career information and the difficulty firms face in appropriating its benefits. Terviö, (2009) shows how markets can under-invest in talent discovery when the returns to information about novice workers are largely external. Empirically, making performance public on online labor markets or via credible third parties causally improves workers’ subsequent outcomes (Pallais,, 2014; Stanton and Thomas,, 2016; Horton,, 2017). Our theory provides conditions under which workers voluntarily seek such transparency by going solo, and conditions under which they rationally prefer opaque employment.
A large empirical literature documents how employers learn about worker productivity over time and the extent to which that learning is public or private (Altonji and Pierret,, 2001). We contribute a distinct mechanism: when firm performance is not disclosed externally, the act of applying for such employment becomes informative to sophisticated employers. This echoes broader insights that markets learn not only from realized outcomes but also from selection and participation decisions. In equilibrium, wages and hiring thresholds internalize the informational selection into opaque jobs versus continued public production, and in our sophisticated pricing case wages are set using the expected applicant pool even when post-hire outputs remain opaque.
Our applications to open-source contributions and freelancing platforms connect to research on reputation systems and public track records (Bar-Isaac and Tadelis,, 2008). In such settings, outside audiences can readily observe some outputs, while many internal tasks remain hidden. The predictions of our model—about switching, absorbing hidden employment for weaker reputations, and wage dynamics across transparency regimes—organize a number of recent empirical facts and suggest new tests, such as how shocks to the visibility of output (e.g., the introduction of new public metrics) shift the incidence of self-employment and the slope of wage–reputation profiles across markets.
Across these strands, our contribution is to endogenize the worker’s choice between public and hidden production, to characterize the absorbing hidden-employment region in which workers rationally remain opaque, and to compare naïve and sophisticated wage-setting regimes that differ in how they map public reputation into wages when some production is deliberately hidden.
3 Model
Time is discrete, $t=0,1,2,\dots$, and the common discount factor is $\delta\in(0,1)$. A single worker has a fixed talent $\theta\in[0,1]$ that is known to the worker and unknown to the outside market. [^6] Formally, one can embed this “single worker” problem in a stationary overlapping-generations environment with a constant exogenous exit probability $\lambda\in(0,1)$ each period. The worker’s Bellman equation then has the continuation term multiplied by $\delta(1-\lambda)$ rather than $\delta$, so all of our results go through with an effective discount factor that incorporates both time preference and expected career length. We adopt this reduced-form formulation in the text for notational simplicity. Firms and outside employers do not observe the worker’s calendar age or total tenure; they only see the history of public outcomes generated under self-employment, which we later summarize by the Beta posterior parameters $(\alpha,\beta)$.
In each period the worker chooses between (i) self-employment ($S$), which yields a binary outcome $y_{t}\in\{0,1\}$ with $\Pr(y_{t}=1\mid\theta)=\theta$ that is publicly observed by the market, and (ii) employment at a firm ($E$), which pays a flat wage $w_{t}$ but keeps the worker’s individual outcome hidden from the outside market. In our baseline, the firm does not disclose individual outcomes externally. [^7] In Section 6 we allow the firm to observe a private signal and discuss renegotiation.
Consumption in period $t$ equals $c_{t}=y_{t}$ under $S$ and $c_{t}=w_{t}$ under $E$ (we normalize the monetary payoff of a success to $1$). The worker has Bernoulli utility $u:\mathbb{R}_{+}\to\mathbb{R}$ that is increasing and concave, with normalization $u(0)=0$ and u(1)=1u(1)=1888This pins units so the expected period utility from self‐employment equals the success probability $\theta$. Any positive affine transformation of $u$ would merely re‐scale wages; all choices are invariant.. Hence the current utility from $S$ is $\theta$ and from $E$ is $u(w_{t})$. Risk aversion (concavity of $u$) matters through comparison with the certain wage under $E$ and through the value of reputation for future wages.
[…]
$\mu(\alpha,\beta)\;=\;\frac{\alpha}{\alpha+\beta}.$
A period of hidden employment leaves $(\alpha,\beta)$ unchanged.
We write $\mu_{t}\equiv\mu(\alpha_{t},\beta_{t})$.
Each period a competitive fringe of short-lived firms posts a flat wage for applicants. Firms cannot condition the wage on contemporaneous individual performance and do not disclose worker-specific outcomes to the market. [^10] Think unverifiable or multi‐task output, team production, or high disclosure costs. Allowing the firm to privately observe a signal and renegotiate does not restore public learning; the absorbing logic only vanishes when the market receives a sufficiently informative public signal from firm employment (Section 6). A firm that hires a worker of talent $\theta$ produces expected revenue $\theta$ (equal to the success probability) and pays wage $w_{t}$. Competition and free entry imply zero expected profits, so the posted wage equals the expected talent of the applicant pool relevant to that firm in that period.
We study two information benchmarks:
  • Naïve firms. Firms ignore the informational content of applying for an opaque firm job. They price off the public state only, posting
    $w^{N}(\alpha,\beta)\;=\;\mu(\alpha,\beta).$ (1)
  • Sophisticated firms. Firms correctly infer that application/acceptance is endogenous and update from the application decision. Let $\Theta^{E}(\alpha,\beta)\subseteq[0,1]$ denote the set of types who apply at state $(\alpha,\beta)$ under the worker’s strategy. Then posted wages satisfy
    $w^{S}(\alpha,\beta)=\mathbb{E}[\theta\mid\theta\in\Theta^{E}(\alpha,\beta);\alpha,\beta],$ (2)
    with the expectation taken under the $\mathrm{Beta}(\alpha,\beta)$ posterior. [^11] With a continuum of firms and applicants, the posted wage is the competitive price for the marginal applicant; with anonymous jobs and free entry, this equals the pool mean talent.
Throughout, we use the term reputation to refer to the public belief about the worker’s talent that is encoded in the posterior distribution and, in particular, in its mean $\mu(\alpha,\beta)$. This public belief is what firms use to set wages in both the naïve and sophisticated cases. In this sense, a worker’s “career concerns” are concerns about how current choices affect future realizations of $\mu(\alpha,\beta)$ and thus future wage offers.
In equilibrium under mild regularity, the worker’s strategy is monotone in $\theta$ and $\Theta^{E}(\alpha,\beta)$ is an interval $0,\hat{\theta}(\alpha,\beta)]$; then $w^{S}$ is the truncated-posterior mean.
At the start of period $t$, the public state $(\alpha_{t},\beta_{t})$ is common knowledge. Firms simultaneously post wages as in (1) or (2) given their inference rule. The worker observes $(\alpha_{t},\beta_{t})$ and his own $\theta$ and chooses $a_{t}\in\{S,E\}$. If $a_{t}=S$, the realization $y_{t}\in\{0,1\}$ is drawn and publicly observed, and the state updates to $(\alpha_{t+1},\beta_{t+1})=(\alpha_{t}+1,\beta_{t})$ if $y_{t}=1$ and $(\alpha_{t},\beta_{t}+1)$ if $y_{t}=0$. If $a_{t}=E$, the worker is paid $w_{t}$ and the public state remains $(\alpha_{t+1},\beta_{t+1})=(\alpha_{t},\beta_{t})$.
Let $V(\theta;\alpha,\beta)$ denote the worker’s value function. For a given wage policy $w(\alpha,\beta)$ (naïve or sophisticated), the Bellman equation is
$V(\theta;\alpha,\beta)$ $=\max\left\{\,U_{S}(\theta;\alpha,\beta),\;U_{E}(\theta;\alpha,\beta)\,\right\},$ (3)
$U_{S}(\theta;\alpha,\beta)$ $=\theta+\delta\Big[\,\theta\,V(\theta;\alpha+1,\beta)+(1-\theta)\,V(\theta;\alpha,\beta+1)\,\Big],$ (4)
$U_{E}(\theta;\alpha,\beta)$ $=u\!\left(w(\alpha,\beta)\right)+\delta\,V(\theta;\alpha,\beta).$ (5)
The term $\theta$ in (4) is current utility $\mathbb{E}[u(y_{t})\mid S,\theta]$ under our normalization $u(1)=1$, $u(0)=0$.
Define
$\Delta(\theta;\alpha,\beta)\equiv U_{S}(\theta;\alpha,\beta)-U_{E}(\theta;\alpha,\beta).$
Under (3)–(5), $\Delta(\theta;\alpha,\beta)$ is strictly increasing in $\theta$: higher $\theta$ raises current expected utility under $S$ and increases the probability of transitioning to the higher-mean state $(\alpha+1,\beta)$. [^12] Formally, $U_{S}$ is strictly increasing in $\theta$, while $U_{E}$ depends on $\theta$ only through the continuation value $V(\theta;\alpha,\beta)$, which is increasing but with slope bounded by $\delta$; the current term in $U_{S}$ delivers the strict single crossing. Hence there exists a unique cutoff $\hat{\theta}(\alpha,\beta)\in[0,1]$ such that types $\theta\leq\hat{\theta}(\alpha,\beta)$ prefer employment and types $\theta>\hat{\theta}(\alpha,\beta)$ prefer self-employment:
$\theta\leq\hat{\theta}(\alpha,\beta)\;\Longleftrightarrow\;U_{E}(\theta;\alpha,\beta)\geq U_{S}(\theta;\alpha,\beta).$ (6)
Monotonicity also implies $\Theta^{E}(\alpha,\beta)=[0,\hat{\theta}(\alpha,\beta)]$.
Definition 1.
[…]
  1. 1.
    Worker optimality. Given $w(\cdot)$ and beliefs, the worker’s policy is (6) and $V$ solves (3)–(5).
  2. 2.
    Firm optimality and zero profits. In the naïve benchmark, $w=w^{N}$ in (1). In the equilibrium case, wages satisfy $w=w^{S}$ in (2) given the implied applicant set $0,\hat{\theta}(\alpha,\beta)]$.
  3. 3.
    Belief consistency. Beliefs update by Bayes’ rule after public outcomes of $S$ (Beta posterior updates) and remain at $(\alpha,\beta)$ after $E$ (no public signal).
In the sophisticated pricing regime, the worker’s cutoff condition (6) and the wage equation (2) jointly pin down the fixed point $(\hat{\theta}(\alpha,\beta),w^{S}(\alpha,\beta))$.
Given the infinite-horizon, overlapping-generations interpretation in Section 3, we focus on stationary Markov equilibria in which the worker’s strategy and wage functions depend only on the current public state $(\alpha,\beta)$ and, in the sophisticated case, on the implied applicant pool. The equilibrium characterized above is stationary in this sense: $(\sigma,w^{N},w^{S})$ do not depend on calendar time. Together with the constant inflow of new workers with prior $(\alpha_{0},\beta_{0})$ and the law of motion for beliefs, this induces a steady-state cross-sectional distribution of public states in the population. Our comparative statics can therefore be interpreted as describing how the steady-state allocation of workers between public and hidden employment responds to changes in primitives such as risk aversion, patience, and performance transparency.
A state–type pair $(\alpha,\beta,\theta)$ is said to be absorbing in employment if, conditional on choosing $E$ at $(\alpha,\beta)$, the worker strictly prefers $E$ again at $(\alpha,\beta)$ next period. Because $E$ generates no public signal, the public state remains $(\alpha,\beta)$, so the cutoff $\hat{\theta}(\alpha,\beta)$ is unchanged. The single-crossing property immediately yields:
Proposition 1.
Fix a public state $(\alpha,\beta)$. For any $\theta\leq\hat{\theta}(\alpha,\beta)$, if the worker chooses $E$ at $(\alpha,\beta)$, then $(\alpha,\beta,\theta)$ is absorbing in employment: the worker chooses $E$ in all subsequent periods. For any $\theta>\hat{\theta}(\alpha,\beta)$, the worker rejects $E$ at $(\alpha,\beta)$ and chooses $S$.
Intuitively, for low-reputation public states $(\alpha,\beta)$ the combination of (i) a low posted wage (especially when firms are sophisticated and price the selected applicant pool) and (ii) risk aversion makes the certain wage comparatively attractive for lower types. Because hidden employment produces no new public information, the public belief $(\alpha,\beta)$ does not improve, and the same comparison repeats in every subsequent period. Proposition 1 therefore implies a simple geometry for the absorbing hidden-employment region: at any public state $(\alpha,\beta)$ the set of types who eventually remain in employment forever is downward closed in talent and can be written as
$\mathcal{A}(\alpha,\beta)\;=\;\left\{\,(\alpha,\beta,\theta):\theta\leq\hat{\theta}(\alpha,\beta)\,\right\},$
with all $\theta\leq\hat{\theta}(\alpha,\beta)$ optimally choosing $E$ and generating no further public signals, while all $\theta>\hat{\theta}(\alpha,\beta)$ continue to self-employ to keep the option value of information alive. Along the path of a given true talent $\theta$, entry into this absorbing region can occur only after a sufficiently adverse sequence of public outcomes drives the public belief $\mu(\alpha,\beta)$ low enough that the insurance value of the flat wage outweighs the option value of another signal. Later, Proposition 5 shows that under sophisticated firms the cutoff satisfies $\hat{\theta}^{S}(\alpha,\beta)\leq\hat{\theta}^{N}(\alpha,\beta)$, so the absorbing region in $(\alpha,\beta,\theta)$-space is weakly smaller when firms fully internalize application-based inference. [^13] Formally, the public state $(\alpha,\beta)$ is a controlled Markov state. Choosing $E$ induces the deterministic transition $(\alpha,\beta)\to(\alpha,\beta)$, so any policy that prescribes $E$ for all $\theta\leq\hat{\theta}(\alpha,\beta)$ is self-confirming at a state where it is optimal.
In the naïve benchmark, wages equal the public mean $\mu(\alpha,\beta)$ regardless of application. In the sophisticated pricing regime, the wage equals the mean of the truncated posterior on $0,\hat{\theta}(\alpha,\beta)]$. The two benchmarks therefore generate different cutoffs: the sophisticated wage $w^{S}$ incorporates the adverse selection induced by (6), which generally lowers $w$ relative to $w^{N}$ and feeds back into the worker’s choice. Section 4 characterizes the fixed point and shows how pricing sophistication shifts $\hat{\theta}(\alpha,\beta)$ and the size of the absorbing region.
Remarks
(i) The Beta prior delivers a two-dimensional sufficient statistic $(\alpha,\beta)$ and clean updating; none of our qualitative results hinge on conjugacy. (ii) Allowing the current employer to observe a private performance signal and renegotiate wages (without public disclosure) does not restore public learning; Section 6 shows that the key absorbing logic persists. (iii) Off-path beliefs about quitting from firm employment can be parameterized by a single pessimism parameter $\kappa\in[0,1]$ without affecting the baseline results; we return to this in Section 4.
4 Equilibrium and characterization
This section characterizes equilibrium policies and wages under (i) a benchmark with naïve pricing and (ii) sophisticated pricing where firms infer from applications. We also establish the existence of an absorbing employment region and provide comparative statics in patience and risk aversion.
4.1 Myopic wage posting
In the benchmark case, the posted wage depends only on the public state: [^14] Competition and free entry deliver zero expected profits. With a positive profit wedge $\pi$, the analysis goes through with $w^{N}=\mu-\pi$, shifting the cutoff monotonically.
$w^{N}(\alpha,\beta)=\mu(\alpha,\beta)\equiv\frac{\alpha}{\alpha+\beta}.$ (7)
If the worker chooses employment, the public state does not change; if she chooses self-employment, the state updates to $(\alpha+1,\beta)$ after a success and to $(\alpha,\beta+1)$ after a failure. Define
$\begin{split}\mu^{+}(\alpha,\beta)\equiv\mu(\alpha+1,\beta)&=\mu+\frac{\beta}{(\alpha+\beta)(\alpha+\beta+1)},\\ \mu^{-}(\alpha,\beta)\equiv\mu(\alpha,\beta+1)&=\mu-\frac{\alpha}{(\alpha+\beta)(\alpha+\beta+1)}.\end{split}$
Let $\Delta(\theta;\alpha,\beta)\equiv U_{S}(\theta;\alpha,\beta)-U_{E}(\theta;\alpha,\beta)$ with $U_{S},U_{E}$ as in (4)–(5). The function $\Delta(\theta;\alpha,\beta)$ is strictly increasing in $\theta$, hence there exists a unique cutoff $\hat{\theta}^{N}(\alpha,\beta)\in[0,1]$ such that types $\theta\leq\hat{\theta}^{N}(\alpha,\beta)$ choose employment and types $\theta>\hat{\theta}^{N}(\alpha,\beta)$ choose self-employment:
$\theta\leq\hat{\theta}^{N}(\alpha,\beta)\quad\Longleftrightarrow\quad U_{E}(\theta;\alpha,\beta)\geq U_{S}(\theta;\alpha,\beta).$ (8)
Proof sketch. $U_{S}$ is strictly increasing in $\theta$ because both the current term and the transition to the more favorable state $(\alpha+1,\beta)$ scale with $\theta$. $U_{E}$ depends on $\theta$ only through the continuation value at the same state $(\alpha,\beta)$, hence has a weaker slope. Therefore $\Delta$ is strictly increasing and crosses zero once. ∎
Because employment leaves the public state unchanged, the cutoff at $(\alpha,\beta)$ is time invariant.
Proposition 2.
Fix $(\alpha,\beta)$. If the worker chooses employment at $(\alpha,\beta)$ and $\theta\leq\hat{\theta}^{N}(\alpha,\beta)$, then she chooses employment forever in all subsequent periods. If $\theta>\hat{\theta}^{N}(\alpha,\beta)$, she rejects employment at $(\alpha,\beta)$ and chooses self-employment.
The absorbing region is therefore
$\mathcal{A}^{N}\equiv\left\{\,(\alpha,\beta,\theta):\theta\leq\hat{\theta}^{N}(\alpha,\beta)\,\right\}.$
The cutoff $\hat{\theta}^{N}(\alpha,\beta)$ depends on (i) the posted wage $\mu(\alpha,\beta)$, (ii) the discount factor $\delta$, and (iii) preferences $u(\cdot)$.
Proposition 3.
(i) $\hat{\theta}^{N}(\alpha,\beta)$ is weakly increasing in $\mu(\alpha,\beta)$: a higher public mean raises the wage and expands the employment region. (ii) $\hat{\theta}^{N}(\alpha,\beta)$ is weakly decreasing in $\delta$: greater patience raises the value of generating public information and shrinks the employment region. (iii) Holding $w^{N}$ fixed, if $u$ becomes more concave in the sense of second-order stochastic dominance (i.e., the worker becomes more risk averse), then $\hat{\theta}^{N}(\alpha,\beta)$ weakly increases. [^15] Risk aversion affects the comparison through $u(w^{N})$ in $U_{E}$. With our normalization $u(1)=1$, $u(0)=0$, the current-period utility from self-employment equals $\theta$, while the utility from the certain wage is $u(\mu)\leq\mu$. A more concave $u$ lowers $u(\mu)$ and therefore makes employment less attractive at the same wage; at the equilibrium cutoff, this translates into a weak increase of $\hat{\theta}^{N}$.
Proof sketch. (i) follows from a monotone shift of $U_{E}$ in $w$; (ii) follows because the marginal value of a public signal (the option value embedded in $U_{S}$) scales with $\delta$ while $U_{E}$ holds the state fixed; (iii) is immediate from the comparison $u(\mu)$ vs. $\theta$ at the cutoff. ∎
As $(\alpha+\beta)$ grows, the length of the worker’s public track record increases and the posterior about talent becomes more precise. A single additional success or failure then moves beliefs by less than at earlier states. Holding the current mean belief $\mu(\alpha,\beta)$ fixed, the one-period option value of an extra public signal is therefore larger when the public record is short and smaller when it is long. Combined with risk aversion, this makes continued self-employment most attractive when the worker has accumulated relatively few public signals (the reputational stakes per signal are high), and makes hidden employment comparatively more attractive once a long track record is already in place. In other words, moving across states in the direction of larger $(\alpha+\beta)$ shifts more types into the absorbing hidden-employment region: at otherwise similar belief states, workers with long public histories are more willing to accept the firm’s insurance and stop generating signals than workers with very short histories.
4.2 Sophisticated wage posting and selection into employment
With sophisticated firms, wages reflect selection into employment: given a cutoff policy $\hat{\theta}(\alpha,\beta)$, the applicant pool at state $(\alpha,\beta)$ is $\Theta^{E}(\alpha,\beta)=[0,\hat{\theta}(\alpha,\beta)]$, and the competitive wage equals the pool mean talent under the $\mathrm{Beta}(\alpha,\beta)$ posterior truncated to this interval:
$w^{S}\!\left(\alpha,\beta;\hat{\theta}\right)\;=\;\mathbb{E}\!\left[\theta\mid\theta\in[0,\hat{\theta}];\;\text{Beta}(\alpha,\beta)\right]\;=\;\frac{B(\hat{\theta};\,\alpha+1,\beta)}{B(\hat{\theta};\,\alpha,\beta)},$ (9)
where $B(c;a,b)\equiv\int_{0}^{c}t^{a-1}(1-t)^{b-1}\,dt$ is the (lower) incomplete beta function.
[…]
$\Delta\!\left(\hat{\theta};\alpha,\beta;\,w^{S}(\alpha,\beta;\hat{\theta})\right)=0,$ (10)
where $\Delta$ is strictly increasing in its first argument.
Lemma 1.
[…]
Proposition 4.
For any $(\alpha,\beta)$ and $\delta\in(0,1)$, there exists a unique cutoff $\hat{\theta}^{S}(\alpha,\beta)\in[0,1]$ solving (10). The corresponding sophisticated wage is $w^{S}(\alpha,\beta;\hat{\theta}^{S})$ from (9). The employment region $\mathcal{A}^{S}\equiv\{(\alpha,\beta,\theta):\theta\leq\hat{\theta}^{S}(\alpha,\beta)\}$ is absorbing.
[…]
Sophisticated pricing internalizes adverse selection: $w^{S}(\alpha,\beta;\hat{\theta})\leq w^{N}(\alpha,\beta)$ with strict inequality whenever $\hat{\theta}<1$. Lower wages tighten the employment condition and shift the cutoff down.
Proposition 5.
For any $(\alpha,\beta)$, $\hat{\theta}^{S}(\alpha,\beta)\leq\hat{\theta}^{N}(\alpha,\beta)$, with a strict inequality when $\hat{\theta}^{N}(\alpha,\beta)<1$.
Proof sketch. For a given state, $w^{S}\leq w^{N}$ because the applicant pool mean does not exceed the unconditional posterior mean. Since $U_{E}$ is increasing in the wage and $\Delta$ is increasing in $\theta$, the indifference point shifts left. ∎
4.3 Comparative statics and testable implications
Increasing $\delta$ raises the option value of generating public signals. In both pricing regimes, $\hat{\theta}(\alpha,\beta)$ is weakly decreasing in $\delta$. [^17] At the sophisticated regime this follows from the envelope property of $V$ and the fact that employment leaves the public state unchanged while self-employment stochastically moves it toward higher $\mu$ with probability $\theta$.
Holding the wage schedule fixed, greater concavity of $u$ weakly increases $\hat{\theta}$ (employment becomes less attractive at a given wage because $u(w)$ falls). With sophisticated pricing, an additional (general-equilibrium) effect operates through composition: when fewer types apply, the applicant pool is stronger, which increases $w^{S}$ and partially offsets the direct effect. Overall, $\hat{\theta}^{S}$ remains weakly increasing in risk aversion under standard regularity.
As $(\alpha+\beta)$ increases, the one-step drift $|\mu^{\pm}-\mu|$ shrinks, reducing the informational benefit of an additional public signal. Hence $\hat{\theta}(\alpha,\beta)$ is weakly increasing in $(\alpha+\beta)$, and the absorbing employment region expands as the public track record becomes longer.
The cutoff $\hat{\theta}(\alpha,\beta)$ links cross-sectional variation in self-employment, wages, and information production. At a given public state $(\alpha,\beta)$, only types with $\theta>\hat{\theta}(\alpha,\beta)$ choose self-employment and continue to generate public signals; lower types choose hidden employment and stop. This implies that occupations or markets in which self-employment or public projects are more prevalent are precisely those in which the cutoff is relatively low and many types reveal themselves, whereas markets with high cutoffs feature both more opaque employment and less public information about talent. The distinction between naïve and sophisticated firms amplifies these patterns. Naïve firms post the public-mean wage $\mu(\alpha,\beta)$ and therefore pay the same wage to all workers with the same public record, regardless of how many low types would be willing to hide. Sophisticated firms anticipate selection into opaque jobs and post a lower wage in states where low types are more likely to accept employment, compressing wages but inducing more self-employment and more public signal production. Cross-sectionally, we therefore expect (i) more self-employment and richer public track records in markets where firms are sophisticated and performance technologies are transparent, and (ii) higher average wages but less public information in markets with naïve pricing and opaque technologies at the same public state.
If the public observes quits from opaque employment, off-path beliefs can be parameterized by $\kappa\in[0,1]$ measuring the pessimism attached to an unexplained quit (Section 6). For a range of $\kappa$, the absorbing region characterized above is unchanged, because beliefs do not update while the worker remains in opaque employment.
Taken together, these comparative statics suggest a simple answer to the question of which labor markets generate the most public information about talent. Markets with more patient and less risk-averse workers, with more transparent performance technologies, and with sophisticated rather than naïve wage setting all have smaller absorbing hidden-employment regions. By Proposition 3, higher patience and better insurance reduce the cutoff $\hat{\theta}(\alpha,\beta)$, while greater risk aversion and a lower reward for public output increase it; by Proposition 5, sophisticated firms choose lower wages than naïve firms at a given public state and thus induce more workers to remain in public production. In such markets, more types experiment with self-employment or public projects, generating richer public track records. Conversely, markets with impatient, highly risk-averse workers, opaque production technologies, and naïve wage setting have larger absorbing regions and slower public learning.
5 Empirical implications
Patience
The discount factor $\delta$ scales the value of future wage improvements that come from generating public signals. At a given public state $(\alpha,\beta)$, let $\Delta(\theta;\alpha,\beta,\delta)$ denote the expected discounted utility difference between choosing $S$ and $E$ for type $\theta$.
By construction, $\Delta$ can be written as a current-period payoff difference plus $\delta$ times the difference in continuation values. The current-period term does not depend on $\delta$, while the continuation term is strictly increasing in both $\theta$ and $\delta$: higher types benefit more from the option value of moving the public belief toward higher $\mu(\alpha,\beta)$, and this option value matters more when the worker is patient. As a result, $\Delta$ has increasing differences in $(\theta,\delta)$, and the unique cutoff $\hat{\theta}(\alpha,\beta;\delta)$ solving $\Delta(\hat{\theta};\alpha,\beta,\delta)=0$ is weakly decreasing in $\delta$.
More patient workers are therefore more willing to forgo the insurance of $E$ in order to refresh their reputation, and absorbing hidden employment is more prevalent when workers are impatient or face short effective horizons.
Risk aversion
Risk aversion makes the certain wage relatively more attractive than the risky Bernoulli payoff from $S$. To see this, consider a one-parameter family of utilities $\{u_{\rho}\}_{\rho}$ on $[0,1]$ with $u_{\rho}(0)=0$ and $u_{\rho}(1)=1$, where a higher $\rho$ indexes greater concavity in the sense of Pratt. At a fixed state $(\alpha,\beta)$ and wage schedule $w(\alpha,\beta)$, the current-period payoff under $E$ is the sure amount $u_{\rho}(w)$, whereas under $S$ it is the expected utility of a Bernoulli payoff with mean $\theta$.
Standard comparative statics of risk aversion imply that, holding the mean fixed, the certainty equivalent of the risky payoff falls as $\rho$ increases. Thus the gain from choosing $S$ over $E$ is pointwise decreasing in $\rho$, and the cutoff $\hat{\theta}(\alpha,\beta;\rho)$ that solves $\Delta(\hat{\theta};\alpha,\beta,\rho)=0$ is weakly increasing in $\rho$. More risk-averse workers are therefore more willing to accept hidden employment at a given public state.
Under sophisticated wage setting, a composition effect reinforces this comparative static: when more low- and mid-types enter $E$, the remaining pool of applicants is stronger, raising the sophisticated wage $w^{S}(\alpha,\beta)$ and further tilting the trade-off toward taking the firm job for marginal types.
Public-track-record length
The length of the public record affects how informative a marginal success or failure is. With a Beta posterior, the public mean is $\mu(\alpha,\beta)=\alpha/(\alpha+\beta)$.
After a success or failure under $S$, the mean updates to
$\begin{split}\mu_{+}(\alpha,\beta)&=\frac{\alpha+1}{\alpha+\beta+1},\\ \mu_{-}(\alpha,\beta)&=\frac{\alpha}{\alpha+\beta+1},\end{split}$
so that
$\begin{split}\mu_{+}(\alpha,\beta)-\mu(\alpha,\beta)&=\frac{\beta}{(\alpha+\beta)(\alpha+\beta+1)},\\ \mu(\alpha,\beta)-\mu_{-}(\alpha,\beta)&=\frac{\alpha}{(\alpha+\beta)(\alpha+\beta+1)}.\end{split}$
For a given mean $\mu$, both expressions shrink like $1/(\alpha+\beta+1)$ as the effective sample size $(\alpha+\beta)$ grows. The option value of an extra public signal therefore declines with the length of the public track record.
Holding wages fixed, $\Delta(\theta;\alpha,\beta)$ shifts down as $(\alpha+\beta)$ increases, so the cutoff $\hat{\theta}(\alpha,\beta)$ is weakly increasing in $(\alpha+\beta)$. In steady state, absorbing hidden employment is thus more prevalent in parts of the labor market where typical workers already have long public histories, while early-career workers with short records have stronger incentives to continue producing public signals.
Transparency premia across occupations
These comparative statics carry clear cross-occupation implications. In occupations with portable public histories—open-source software, contests, online freelancing platforms, or fields with visible publication and citation records—workers can more easily accumulate long public track records.
Our model predicts (i) a higher prevalence of self-employment or public projects among workers with strong reputations in such occupations, because a lower cutoff $\hat{\theta}(\alpha,\beta)$ makes public production attractive for many types; (ii) sharper wage increases after publicly observed successes and steeper penalties after public failures, since wages respond more to public signals when they are informative; and (iii) more frequent transitions into opaque employment following runs of poor public realizations, especially for mid-ability workers whose reputations are pushed toward the absorbing hidden-employment region.
These patterns align with evidence that public ratings and track records on digital platforms and innovation ecosystems have first-order effects on hiring and earnings, and that workers respond by reallocating effort toward more visible tasks when the returns to transparency are high. [^18] See, for example, empirical work on online labor markets and open-source contributions that documents how public performance measures shape demand, wages, and workers’ portfolio choices.
Pricing sophistication and wage dispersion
Sophisticated wage setting makes the adverse selection in applicant pools explicit. In the naïve benchmark, wages equal the public posterior mean $w^{N}(\alpha,\beta)=\mu(\alpha,\beta)$ regardless of who is expected to apply.
Under sophisticated pricing, wages equal the mean talent of the implied applicant pool $0,\hat{\theta}(\alpha,\beta)]$, which is strictly below $\mu(\alpha,\beta)$ whenever $\hat{\theta}(\alpha,\beta)<1$. Thus at any given public state, sophisticated wages $w^{S}(\alpha,\beta)$ are lower on average than naïve benchmark wages, reflecting a familiar adverse-selection discount. This compresses wage dispersion after public successes and failures: when a positive public shock moves the state from $(\alpha,\beta)$ to $(\alpha+1,\beta)$, naïve wages jump by $\mu_{+}-\mu$, whereas sophisticated wages move by the change in the truncated-posterior mean.
Because sophisticated wages already incorporate who is expected to apply, the incremental impact of a public success on wages is smaller. Relative to naïve environments in which HR ignores selection into opaque jobs, sophisticated pricing therefore implies (i) lower average wages for otherwise identical public states, (ii) fewer mid-type applicants for opaque jobs (some mid-ability workers prefer to keep generating public signals to avoid the adverse-selection discount), and (iii) stronger sorting of workers into public versus hidden production based on outside reputations.
Empirical probes
The model suggests a set of empirical probes that map directly into the comparative statics above.
First, event-study designs can compare wage paths after public successes and failures for workers who remain in self-employment versus those who move into opaque employment, conditioning on the public state $(\alpha,\beta)$ at the time of the shock. The theory predicts larger wage responses and slower entry into opaque jobs after positive shocks in transparent, naïve-pricing environments, and more muted wage dispersion when pricing is sophisticated.
Second, cross-occupation comparisons can relate the share and timing of opaque spells to measures of transparency and public-track-record length: occupations with more portable public histories should exhibit earlier switching into hidden employment for weaker reputations and stronger persistence in opaque spells once workers enter the absorbing region.
Third, within-platform variation in “pricing sophistication”—for instance, across employers with different screening technologies or across platforms with more versus less curated rating systems—can be used to test whether mid-quantile workers are more likely to keep generating public signals when they anticipate stronger adverse-selection discounts in opaque jobs.
Together, these tests speak to how selection into self-employment versus hidden jobs and the sophistication of wage setting jointly determine both wage dynamics and the amount of public information that labor markets generate about talent.
6 Discussion and extensions
This section discusses how our mechanisms interact with (i) worker learning about own type, (ii) partial public observability inside firms, and (iii) simple screening menus. The baseline insights—a cutoff policy with an absorbing hidden-employment region and selection into public production—are robust.
6.1 Worker learning about own type
In the baseline, $\theta$ is fixed and privately known. Here the worker starts with a private prior $H_{0}$ and updates after each period using privately observed outcomes; public outcomes of $S$ update both $H_{t}$ and the market’s $F_{t}$, while $E$ updates only $H_{t}$. Let $m_{t}\equiv\mathbb{E}[\theta\mid\text{worker's private history}]$ and $\mu_{t}\equiv\mathbb{E}[\theta\mid\text{public history}]$.
Remark 1.
Fix a public state $(\alpha,\beta)$ (summarized by $\mu_{t}$). For any wage schedule and $\delta\in(0,1)$, there exists a unique cutoff $\hat{m}(\alpha,\beta)$ such that the worker chooses $E$ iff $m_{t}\leq\hat{m}(\alpha,\beta)$; the map $m\mapsto U_{S}(m;\alpha,\beta)-U_{E}(m;\alpha,\beta)$ is strictly increasing.
The absorbing logic extends: if $E$ is optimal at $(\mu_{t},m_{t})$, then the public state remains $(\alpha,\beta)$ and the same comparison repeats. [^19] Private updating under $E$ can move $m_{t+1}$, but with flat wages independent of private realizations, there is a neighborhood of private beliefs for which absorption persists. Public learning is restored only if the market sees an informative signal from $E$. Private learning under $E$ may move $m_{t+1}$ but does not overturn absorption locally with flat wages. Under sophisticated posting (application inference), the cutoff $\hat{m}(\alpha,\beta)$ shifts down due to the adverse-selection discount.
6.2 Partial public observability inside firms
So far, employment at the firm $E$ has been fully opaque to the outside market: individual performance affects wages only through its effect on the worker’s continuation value inside the firm. In many settings, however, firm jobs generate some coarse public signals—for example, occasional client feedback, standardized performance ratings, or sporadic disclosure of individual outcomes. To capture this, suppose that under $E$ the market observes a binary signal $z_{t}\in\{0,1\}$ with informativeness $\phi\in[0,1]$:
$\Pr(z_{t}=1\mid\theta,E)\;=\;\phi\,\theta+(1-\phi)\,\bar{z},$
where $\bar{z}\in(0,1)$ is a baseline success probability independent of $\theta$. When $\phi=0$, the signal is uninformative about talent and the extension collapses to the baseline model in which $E$ generates no public information. As $\phi$ increases, $z_{t}$ becomes more informative about $\theta$ in the sense of Blackwell: a larger share of the signal is driven by the same $\theta$-dependent component that governs public output under $S$. In the limit, as $\phi\to 1$, $E$ and $S$ generate nearly parity information about talent.
This partial observability changes the worker’s dynamic trade-off in two ways. First, for high-$\theta$ types, more informative in-firm signals make $E$ more attractive: choosing $E$ no longer freezes the public record, so a talented worker can rely on occasional successes under $E$ to sustain or improve her reputation. Second, for low-$\theta$ types, more informative in-firm signals make $E$ less attractive: opacity is less effective as a shield and the benefit of retreating into hidden employment is reduced.
Consequently, the absorbing hidden-employment region—the set of public states $(\alpha,\beta)$ at which it is optimal for all types to choose $E$ forever and stop actively managing their public record—shrinks as $\phi$ increases. For sufficiently low values of $\phi$, the region remains non-empty: if in-firm signals are very noisy, hidden employment is still effectively opaque and the analysis from Sections 45 carries through. As $\phi$ approaches one, the region eventually disappears: once $E$ and $S$ generate almost equally informative public records, there is no public state at which all types strictly prefer to remain in $E$ forever purely for informational reasons.
The following result formalizes this comparative static.
Proposition 6.
Consider the extension with in-firm signals $z_{t}$ as above and denote by $\mathcal{A}(\phi)\subseteq\{(\alpha,\beta)\}$ the absorbing hidden-employment region under informativeness parameter $\phi$, i.e., the set of public states $(\alpha,\beta)$ at which it is optimal for all $\theta\in[0,1]$ to choose $E$ forever. Then:
  1. 1.
    At $\phi=0$, $\mathcal{A}(0)$ coincides with the absorbing region characterized in Proposition 1 and is non-empty.
  2. 2.
    For $\phi$ in a neighbourhood of zero, $\mathcal{A}(\phi)$ remains non-empty.
  3. 3.
    If at $\phi=1$ the public signals generated under $E$ and $S$ are observationally equivalent (i.e., induce the same conditional distribution of public histories for every $\theta$), then $\mathcal{A}(\phi)$ shrinks as $\phi$ increases and is empty for all $\phi$ sufficiently close to one.
Proof sketch.
At $\phi=0$ the extension reduces to the baseline model, so $\mathcal{A}(0)$ is non-empty by Proposition 1. For $\phi$ in a neighbourhood of zero, continuity of the worker’s value function and of best responses in the signal distribution implies that the set of states at which $E$ strictly dominates $S$ for all types deforms only slightly. Hence, $\mathcal{A}(\phi)$ remains non-empty for all $\phi\in[0,\phi_{0})$ for some $\phi_{0}>0$.
As $\phi$ increases, the signal under $E$ becomes more informative about $\theta$ in the sense of Blackwell, and the transition kernel for public beliefs under $E$ converges to that under $S$. Under the maintained assumption that at $\phi=1$ the distribution of public histories is the same under $E$ and $S$ for any fixed $\theta$, the only remaining difference between the two actions is their one-period payoff streams.
For any public state $(\alpha,\beta)$ there are then types $\theta$ who strictly prefer $S$ (e.g., high types for whom the additional risk under $S$ is rewarded) and types who weakly prefer $E$. Thus, there is no state at which all $\theta$ strictly prefer $E$ forever and $\mathcal{A}(1)$ is empty. Upper hemicontinuity of $\mathcal{A}(\cdot)$ in $\phi$ then implies that $\mathcal{A}(\phi)=\emptyset$ for all $\phi$ sufficiently close to one. ∎
Proposition 6 shows that the absorbing hidden-employment region is robust to modest in-firm transparency but is not a knife-edge artifact of complete opacity. As long as in-firm signals remain sufficiently noisy, there are public states at which all types optimally retreat into hidden employment and allow the public record to stagnate. As in-firm performance becomes clearer to outside observers, this region shrinks and eventually vanishes: opacity inside firms ceases to be a distinct reputational option, and the worker’s problem increasingly resembles choosing between two jobs that generate comparable public records.
This extension connects our mechanism to literatures on partial transparency and signal design inside organizations, where the informativeness of performance measures can be chosen or regulated, and highlights that modest increases in in-firm observability need not overturn the central absorbing-region logic.
6.3 Simple screening menus
So far, employment at the firm $E$ has been represented by a single contract: a flat wage and full opacity to the outside market. Many real jobs, however, combine a fixed wage with a choice over how much of the worker’s output is made visible outside the firm (e.g. more client-facing work, more external presentations, or more disclosure of individual performance metrics). To capture this in reduced form, suppose firms can post simple menus of the form $(w,p)$, where $w$ is the wage and $p\in[0,1]$ is the intensity of public-facing tasks or disclosure in $E$. A higher $p$ means that a larger fraction of the worker’s performance under $E$ is transformed into public signals (in the same spirit as the partial observability parameter $\phi$ in the previous subsection), so that $p=0$ corresponds to a fully opaque firm job and $p=1$ (to first order) to a job that generates almost the same public record as $S$.
For a given menu $(w,p)$ and public state $(\alpha,\beta)$, the worker’s expected utility under $E$ trades off three effects. First, higher $w$ raises current consumption and provides insurance against output risk. Second, higher $p$ makes employment more like self-employment from a reputational point of view: the public record continues to evolve under $E$, which is attractive for high-$\theta$ types and unattractive for low-$\theta$ types. Third, the option value of switching back to self-employment in the future depends on how the public state $(\alpha,\beta)$ evolves under $(w,p)$. For fixed $p$, there is single crossing in $\theta$ between $E$ and $S$ as in the baseline model: types above a cutoff prefer to bear the risk of $S$ in exchange for more informative public signals, while types below the cutoff prefer the firm’s insurance. For fixed $w$, there is also a natural single-crossing structure in $p$: high-$\theta$ types value being observed (higher $p$), low-$\theta$ types value opacity (lower $p$).
Under competitive posting and zero profits, wages must adjust to the menu’s informativeness. When a firm increases $p$ in order to make a contract more attractive to high types, the implied applicant pool becomes more positively selected, but the contract also exposes workers to more reputational risk. Zero profits then require lowering $w$ to keep expected output net of wages at zero. This has two implications. First, menus cannot generate rich separation within employment over and above the self-employment margin: any high-$p$ contract that succeeds in attracting high types must offer a wage low enough that some of those types are again better off choosing $S$, which undermines the intended screening. Second, the main margin remains entry versus self-employment: menus tilt the location of the cutoff $\hat{\theta}(\alpha,\beta)$ by making $E$ more or less attractive at the boundary, but they do not create additional disjoint regions in $(\alpha,\beta,\theta)$ space where different types sort into qualitatively different forms of employment inside the firm.
The upshot is that simple screening menus are unlikely to overturn the absorbing hidden-employment region. For sufficiently low $p$ (or for menus in which the high-$p$ option is not too attractive), the logic of Proposition 1 continues to apply: there remain public states $(\alpha,\beta)$ at which all types strictly prefer firm employment over self-employment, and once a worker enters that region and chooses $E$, the public record ceases to be actively managed. Allowing menus shifts the cutoff and thus the size of the absorbing region—for example, making transparent menus more widely available can shrink the region when workers are sufficiently high types—but does not eliminate the basic absorption logic. A full characterization of optimal menus and their interaction with dynamic reputation is beyond the scope of this paper; here the main lesson is that simple competitive screening devices at the firm level tilt the entry margin rather than provide a complete substitute for self-employment as a tool for reputation management.
6.4 Hidden effort and work–shirk dynamics
So far, self-employment affects only the observability of output: conditional on choosing $S$, success arrives with probability $\theta$ and failure with probability $1-\theta$. A natural extension is to allow the worker to choose a hidden effort level in self-employment, with effort affecting the success probability. For instance, if the worker exerts effort $e\in\{0,1\}$ at a private cost $c>0$, we might have
$\Pr(y_{t}=1\mid\theta,e=1)=\theta\quad\text{and}\quad\Pr(y_{t}=1\mid\theta,e=0)=\phi\,\theta$
for some $\phi\in(0,1)$; exerting effort raises the chance of success, but effort is not observed by the market. This introduces a moral-hazard margin on top of the observability choice.
The qualitative incentives are straightforward. Under $E$, individual output is hidden and wages do not depend on performance in our baseline, so career concerns provide little reason to exert costly effort: a worker who has retreated into hidden employment mainly values the insurance component and, absent additional frictions, will tend to shirk. Under $S$, by contrast, a success is a public signal that improves future wages, while a failure is a public setback. The value of effort in self-employment is therefore tied to how much a marginal success or failure moves the posterior about talent, which in turn depends on the current public state $(\alpha,\beta)$.
This suggests a non-monotone “work–shirk” pattern in reputation space that is closely related to Board and Meyer-ter Vehn, (2013). When the public record is short and beliefs are still diffuse, a single success can substantially raise the posterior mean, so the reputational return to effort in $S$ is high; workers with moderate or promising reputations have strong incentives to work rather than shirk. When the public record is very long and beliefs are precise, each additional success or failure moves $\mu(\alpha,\beta)$ only slightly, so the marginal reputational benefit of effort is smaller; depending on $c$, some high-reputation workers may optimally coast on their existing reputation. At very low reputations, the chance that even a worked-for success would materially improve future wages can be small, so discouraged workers may shirk as well. In short, allowing hidden effort would generate familiar work–shirk dynamics in which effort is most valuable at intermediate reputation levels, echoing the comparative statics of dynamic reputation models under different information structures.
Adding hidden effort would also reinforce the logic of the absorbing region. Once a worker enters the absorbing hidden-employment region, there is little reputational gain from exerting effort in $E$, so both public information production and effort provision are low. By contrast, for types near the cutoff $\hat{\theta}(\alpha,\beta)$, the prospect of escaping absorption by producing a public success can make effort in $S$ particularly attractive. A full solution of the extended model is beyond our scope here, but the discussion above suggests that our main comparative statics—about how risk exposure, transparency, and wage-setting affect the absorbing region and selection into public production—are robust to introducing a hidden effort choice. [^20] We thank an anonymous referee for suggesting that we discuss hidden effort and its connection to work–shirk dynamics in dynamic reputation models, such as Board and Meyer-ter Vehn, (2013).
6.5 Policy implications: taxes and subsidies for public work
The model highlights a simple information externality. When the worker chooses self-employment $S$, the public signal generated in that period not only affects the worker’s own future wages but also improves the allocation of talent in the wider labor market: better information about ability helps firms match workers to jobs more efficiently. By contrast, when the worker chooses hidden employment $E$, individual outcomes remain opaque and the public belief does not move. From a social point of view, workers privately under-produce public signals relative to a planner who internalizes this informational benefit, especially when risk aversion and incomplete insurance make $E$ attractive.
A natural policy lever is to change the relative payoff of $S$ and $E$ through taxes or subsidies on activities that generate public records. Suppose a proportional tax $\tau\in[0,1)$ is levied on self-employment income, so that a success under $S$ yields $(1-\tau)$ instead of $1$. Holding beliefs fixed, this reduces the current-period return to $S$ and raises the talent cutoff $\hat{\theta}(\alpha,\beta)$ in the worker’s indifference condition. In terms of the absorbing hidden-employment region
$\mathcal{A}(\alpha,\beta)=\left\{(\alpha,\beta,\theta):\theta\leq\hat{\theta}(\alpha,\beta)\right\},$
a higher tax expands $\mathcal{A}(\alpha,\beta)$: some types who would have chosen $S$ at a given public state now optimally remain in $E$ and stop generating public information. A symmetric argument applies to a subsidy (a negative $\tau$): by increasing the net payoff from $S$, a subsidy lowers $\hat{\theta}(\alpha,\beta)$, shrinks the absorbing region and induces more experimentation and public signal production. This comparative static is directly analogous to the effects of risk aversion, patience, and wage policies characterized in Proposition 3: policies that effectively make $S$ less attractive (more risky, less remunerative) raise the cutoff and reduce information revelation; policies that make $S$ more attractive do the opposite.
In practice, a planner would trade off the informational benefits of encouraging $S$ against the costs of additional risk and potential misallocation when low-ability workers also experiment. But the qualitative prescription is clear: if the goal is to maximize public information about talent—so that high-$\theta$ workers are identified and deployed efficiently sooner—taxes on activities that generate portable public records (freelancing, open-source contributions, public performance metrics) are counterproductive, while subsidies or insurance for such activities can be welfare-improving. This perspective offers one rationale for policies such as tax breaks for early-career entrepreneurs or support for platforms that document individual performance, and cautions against fiscal measures that inadvertently widen the absorbing region by discouraging public work.
The same logic applies outside labor markets. One can reinterpret the worker as a firm with a privately known product quality and the public signal as a noisy review or rating. Choosing $S$ is then analogous to offering free trials or pay-what-you-want deals that generate public feedback, while choosing $E$ corresponds to selling only through opaque channels where individual customer experiences are not easily observed or shared. High-quality firms are more willing to bear demand risk in order to build a public reputation, whereas low-quality firms prefer the insurance value of opaque pricing or bundled sales.
The absorbing hidden-employment region in our baseline becomes an absorbing hidden-sales region in which firms profitably remain opaque and stop generating public reviews. Comparative statics carry over: environments with better risk sharing, more patient owners, or more transparent technologies induce more firms to experiment with free trials and other reputation-building devices, increasing the public information available about product quality.
7 Conclusion
This paper studies career concerns when some production is publicly observable and some is opaque. A risk-averse worker chooses each period between self-employment, which generates public signals about talent, and hidden firm employment, which pays a flat wage and produces no public information. Firms post wages either by benchmarking myopically on the public state or, in equilibrium, by inferring from applications and pricing the selected applicant pool. The worker’s problem exhibits single crossing, so behavior is governed by a cutoff policy in the public state.
Three forces organize the results. First, an insurance–information trade-off: hidden employment offers insurance but suppresses public learning and future outside options. Second, selection: at any public state there is a unique talent cutoff separating workers who continue to self-employ from those who choose hidden employment. Third, pricing sophistication: in the sophisticated regime, wages are set in anticipation of who will apply, so relative to the myopic benchmark they incorporate an adverse-selection discount. Together these forces generate an absorbing hidden-employment region and a pattern in which higher-talent workers select into public production while lower-talent workers select into hidden jobs.
We deliver a simple characterization under both posting regimes, show how the size of the absorbing region responds to risk aversion, patience, and transparency, and show that sophisticated pricing compresses wages, shrinks the absorbing region, and tightens self-selection by shifting the cutoff downward at each public state.
The framework yields transparent empirical implications. Early in a worker’s public track record—when a single public outcome moves beliefs more— workers are more willing to generate public information; once a long public record is in place and additional signals move beliefs less, absorbing hidden employment becomes more prevalent. After public successes (failures), the benchmark predicts wide wage dispersion, whereas sophisticated posting corrects for winner’s-curse effects through application-based inference, compressing the success–failure gap. Across occupations, environments with portable public histories (e.g., open portfolios, platform ratings, contest leaderboards) should feature stronger selection into public production by higher-ability workers and sharper wage responses to public outcomes. [^21] Examples include code and data science (open repositories), academic research (working papers and preprints), design (public portfolios), and online freelancing (platform ratings).
We showed that the main insights are robust. Allowing the worker to learn about own type preserves the cutoff structure (in private beliefs) and the absorbing logic because hidden employment freezes the public state. Introducing partial public observability inside firms shrinks, but does not eliminate, the absorbing region unless internal signals become nearly as informative as public production. Letting firms offer simple $(w,p)$ menus shifts the cutoff but does not overturn the central margin: the entry decision mediates the trade-off between insurance and reputation.
Two design lessons follow. First, organizational transparency is a lever: policies that selectively externalize performance (client-visible tasks, standardized public metrics) can substitute for self-employment as a reputational technology, [^22] For instance, making some client KPIs or review summaries public, adopting external code review, or giving structured, verifiable titles that map to outside‐market signals. reducing the absorbing hidden-employment region and speeding market learning. Second, marketplace curation (platform labels, verified histories) changes the relative returns to public signals and therefore the equilibrium mix of public versus hidden production.
Naturally, the model is stylized. We abstract from multi-tasking, multi-dimensional ability, strategic project choice, and richer long-term contracts, and we treat firms as short-lived competitors rather than repeat principals. Extending the analysis to relational contracts, internal renegotiation with private performance signals, or general priors beyond the Beta family would sharpen the link to personnel and organizational economics. On the empirical side, the predictions invite structural or event-study tests on settings with observable public tracks and opaque firm spells (e.g., platform markets, open-source ecosystems, or occupations with client-visible KPIs).
By tying a classic career-concerns logic to environments with coexisting public and hidden performance, the paper provides a compact theory of how selection into self-employment versus hidden jobs and the sophistication of wage setting jointly determine when employment becomes absorbing, how wages reflect reputation, and how much public information about talent different labor markets generate.
Our analysis is also closely related to recent work on dynamic reputation with endogenous records and effort. The stationary, overlapping-generations formulation we adopt follows the spirit of Pei, (2025), who shows how record design and long lifetimes shape incentives in repeated interactions; here, the choice between self-employment and hidden employment plays a similar role in shaping the worker’s public record. The hidden-effort extension discussed above links our selection mechanism to the work–shirk dynamics, highlighting that our main comparative statics about transparency and risk exposure continue to matter once effort provision is added to the observability choice.
Appendix A Proofs
Proof of Lemma 1.
With
$U_{S}(\theta;\alpha,\beta)=\theta+\delta\{\theta V(\theta;\alpha{+}1,\beta)+(1{-}\theta)V(\theta;\alpha,\beta{+}1)\}$
and $U_{E}(\theta;\alpha,\beta)=u(w)+\delta V(\theta;\alpha,\beta)$, the current term implies $\partial U_{S}/\partial\theta>0$; the transition also shifts weight toward $(\alpha{+}1,\beta)$, where $V$ is (weakly) higher by value monotonicity. By contrast, $U_{E}$ depends on $\theta$ only through $V(\theta;\alpha,\beta)$ (same state), hence $dU_{E}/d\theta$ is weakly smaller. Therefore $\Delta(\theta;\cdot,w)$ is strictly increasing and continuous in $\theta$.
$U_{E}$ is strictly increasing and continuous in $w$, while $U_{S}$ is $w$-independent. Hence $\Delta$ is strictly decreasing and continuous in $w$.
[…]
Proof of Proposition 1.
Fix a public state $(\alpha,\beta)$ in the absorbing employment region and let $w(\alpha,\beta)$ be the associated wage. For each type $\theta$ let $\Delta(\theta;\alpha,\beta,w(\alpha,\beta))$ denote the expected discounted utility difference between choosing $S$ and choosing $E$ when the current public state is $(\alpha,\beta)$, the wage is $w(\alpha,\beta)$, and future play is given by the equilibrium strategy. By definition of the cutoff $\hat{\theta}(\alpha,\beta)$ we have
$\Delta(\theta;\alpha,\beta,w(\alpha,\beta))\;\leq\;0\quad\text{for all }\theta\leq\hat{\theta}(\alpha,\beta),$
that is, at $(\alpha,\beta)$ these types weakly prefer $E$ to $S$.
Under $E$ the public state is a controlled Markov state with deterministic transition $(\alpha,\beta)\mapsto(\alpha,\beta)$: if the worker chooses $E$ today, the public posterior remains $(\alpha,\beta)$ tomorrow and the firm posts the same wage $w(\alpha,\beta)$ again. Thus, if a type $\theta\leq\hat{\theta}(\alpha,\beta)$ chooses $E$ in the current period, the decision problem he faces next period is identical to the one he faces today: the public state and wage are unchanged and the equilibrium continuation strategy profile is the same.
Let $V_{t}(\alpha,\beta,\theta)$ denote the worker’s continuation value at the start of period $t$, and suppose that at some date $t_{0}$ the public state is $(\alpha,\beta)$ and $\theta\leq\hat{\theta}(\alpha,\beta)$. Consider the stationary strategy that prescribes $E$ in every period $t\geq t_{0}$. Under this strategy the public state remains $(\alpha,\beta)$ forever, so the Bellman equation at $t_{0}$ can be written as
$V_{t_{0}}(\alpha,\beta,\theta)=\max\{U_{S}(\alpha,\beta,\theta),\,U_{E}(\alpha,\beta,\theta)\},$
where $U_{S}$ and $U_{E}$ denote the current-period payoff plus discounted continuation value from playing $S$ and $E$, respectively, given the equilibrium continuation strategy. By construction, $U_{S}(\alpha,\beta,\theta)-U_{E}(\alpha,\beta,\theta)=\Delta(\theta;\alpha,\beta,w(\alpha,\beta))\leq 0$, so the maximum is attained at $E$:
$V_{t_{0}}(\alpha,\beta,\theta)=U_{E}(\alpha,\beta,\theta).$
Exactly the same argument applies at $t_{0}+1,t_{0}+2,\ldots$ because the public state and wage remain $(\alpha,\beta)$ and the continuation strategy is stationary. Hence at every date $t\geq t_{0}$ there is no profitable one-period deviation from the strategy that prescribes $E$.
By the one-step deviation principle for stationary dynamic programming problems, the strategy that chooses $E$ in every period $t\geq t_{0}$ is sequentially optimal for types $\theta\leq\hat{\theta}(\alpha,\beta)$ starting from $(\alpha,\beta)$. In particular, once such a type enters the absorbing employment region and chooses $E$, he will optimally continue to choose $E$ forever and never produce public signals again, so the region is absorbing. ∎
Proof of Proposition 4.
For any $w\in(0,1)$, $\Delta(0;\cdot,w)<0$ (since $U_{E}(0)=u(w)>0=U_{S}(0)$) and $\Delta(1;\cdot,w)>0$ (since $U_{S}(1)=1+\delta V(\cdot)\!>\!u(w)+\delta V(\cdot)=U_{E}(1)$).
By continuity and strict single crossing in $\theta$ there exists a unique $\phi(w)\in(0,1)$ with $\Delta(\phi(w);\cdot,w)=0$, and $\phi(w)$ is strictly decreasing in $w$ by the implicit function theorem and Lemma 1(2).
[…]
Proof of Proposition 5.
Fix a public state $(\alpha,\beta)$ and write $\Delta(\theta,w)\equiv\Delta(\theta;\alpha,\beta,w)$ for the expected discounted utility difference between $S$ and $E$ when the current wage is $w$. By Lemma 1, for any fixed wage $w$ the function $\theta\mapsto\Delta(\theta,w)$ is strictly increasing and continuous on $[0,1]$, and for any fixed type $\theta$ it is strictly decreasing and continuous in $w$.
Let $w^{N}(\alpha,\beta)$ denote the naïve wage and $\hat{\theta}^{N}(\alpha,\beta)$ the corresponding cutoff, so that
$\Delta\big(\hat{\theta}^{N}(\alpha,\beta),\,w^{N}(\alpha,\beta)\big)=0.$
Under sophisticated pricing, the equilibrium wage at $(\alpha,\beta)$ is $w^{S}(\alpha,\beta;\hat{\theta}^{S}(\alpha,\beta))$, where $\hat{\theta}^{S}(\alpha,\beta)$ is the unique cutoff solving $\Delta(\hat{\theta}^{S}(\alpha,\beta),\,w^{S}(\alpha,\beta;\hat{\theta}^{S}(\alpha,\beta)))=0$. Because the sophisticated wage equals the mean of the truncated posterior on $0,\hat{\theta}^{S}(\alpha,\beta)]$, while the naïve wage equals the unconditional posterior mean, we have
$w^{S}(\alpha,\beta;\hat{\theta}^{S}(\alpha,\beta))\;\leq\;w^{N}(\alpha,\beta),$
with strict inequality whenever $\hat{\theta}^{S}(\alpha,\beta)<1$ (truncating a non-degenerate upper tail lowers the mean).
Since $\Delta$ is strictly decreasing in $w$, evaluating at $\theta=\hat{\theta}^{N}(\alpha,\beta)$ gives
$\Delta\big(\hat{\theta}^{N}(\alpha,\beta),\,w^{S}(\alpha,\beta;\hat{\theta}^{S}(\alpha,\beta))\big)\;\geq\;\Delta\big(\hat{\theta}^{N}(\alpha,\beta),\,w^{N}(\alpha,\beta)\big)=0,$
with strict inequality whenever $w^{S}(\alpha,\beta;\hat{\theta}^{S}(\alpha,\beta))<w^{N}(\alpha,\beta)$.
Because $\theta\mapsto\Delta(\theta,w^{S}(\alpha,\beta;\hat{\theta}^{S}(\alpha,\beta)))$ is strictly increasing and crosses zero exactly once, the unique solution $\hat{\theta}^{S}(\alpha,\beta)$ to $\Delta(\theta,w^{S}(\alpha,\beta;\hat{\theta}^{S}(\alpha,\beta)))=0$ must satisfy
$\hat{\theta}^{S}(\alpha,\beta)\;\leq\;\hat{\theta}^{N}(\alpha,\beta),$
with strict inequality whenever $w^{S}(\alpha,\beta;\hat{\theta}^{S}(\alpha,\beta))<w^{N}(\alpha,\beta)$, i.e. whenever $\hat{\theta}^{N}(\alpha,\beta)<1$ and selection into employment is non-trivial. ∎
Proof of Proposition 6.
For each $\phi\in[0,1]$, consider the stationary equilibrium of the extended game in which the in-firm signal under $E$ has distribution $\Pr(z_{t}=1\mid\theta,E)=\phi\theta+(1-\phi)\bar{z}$, while the public signal under $S$ is as in the baseline model. Let $V^{\phi}(\alpha,\beta,\theta)$ denote the worker’s equilibrium value function at public state $(\alpha,\beta)$ and talent $\theta$, and let
$\Delta^{\phi}(\theta;\alpha,\beta)\;\equiv\;\mathbb{E}\big[U^{\phi}\mid S,\theta,(\alpha,\beta)\big]-\mathbb{E}\big[U^{\phi}\mid E,\theta,(\alpha,\beta)\big]$
denote the equilibrium expected discounted utility difference between choosing $S$ and $E$ at $(\alpha,\beta)$ for type $\theta$, where expectations are taken over current payoffs and future public signals given the equilibrium strategies and beliefs. By definition, the absorbing hidden-employment region $\mathcal{A}(\phi)$ is the set of public states $(\alpha,\beta)$ at which $\Delta^{\phi}(\theta;\alpha,\beta)\leq 0$ for all $\theta\in[0,1]$ and strictly for some $\theta$ (i.e., $E$ is strictly preferred to $S$ for at least one type and weakly preferred for all types).
(i) $\mathcal{A}(0)$ coincides with the baseline absorbing region and is non-empty. When $\phi=0$, the signal $z_{t}$ is independent of $\theta$. In this case the extended game reduces exactly to the baseline model in which $E$ generates no public information: the transition kernel for public beliefs under $E$ is the identity mapping, and the worker’s Bellman equation and equilibrium conditions coincide with those in Sections 3 and 4. Hence $\mathcal{A}(0)$ coincides with the absorbing hidden-employment region characterized in Proposition 1, which is non-empty by that result.
(ii) $\mathcal{A}(\phi)$ remains non-empty for $\phi$ in a neighbourhood of zero. Fix $(\alpha^{\ast},\beta^{\ast})\in\mathcal{A}(0)$ such that $\Delta^{0}(\theta;\alpha^{\ast},\beta^{\ast})<0$ for all $\theta\in[0,1]$; such a state exists because Proposition 1 implies that in the baseline there is a region of public states at which hidden employment strictly dominates self-employment for all types in a non-degenerate interval of talents. Since $0,1]\times\{(\alpha^{\ast},\beta^{\ast})\}$ is compact and $\Delta^{0}(\cdot;\alpha^{\ast},\beta^{\ast})$ is continuous in $\theta$, there exists $\varepsilon>0$ such that
$\Delta^{0}(\theta;\alpha^{\ast},\beta^{\ast})\;\leq\;-\varepsilon\quad\text{for all }\theta\in[0,1].$
Next note that, for each fixed $(\theta,\alpha,\beta)$, the current-period payoffs under $S$ and $E$ and the transition probabilities for public beliefs depend continuously on $\phi$. The Bellman operator for the worker’s problem is a contraction with respect to the sup norm and depends continuously on $\phi$ in the sense of Berge’s maximum theorem. It follows that the unique fixed point $V^{\phi}$ is continuous in $\phi$, and hence that $\Delta^{\phi}(\theta;\alpha,\beta)$ is continuous in $(\theta,\alpha,\beta,\phi)$. In particular, $\Delta^{\phi}(\theta;\alpha^{\ast},\beta^{\ast})$ converges uniformly to $\Delta^{0}(\theta;\alpha^{\ast},\beta^{\ast})$ in $\theta$ as $\phi\to 0$. Therefore there exists $\phi_{0}>0$ such that for all $\phi\in[0,\phi_{0})$,
$\sup_{\theta\in[0,1]}\Delta^{\phi}(\theta;\alpha^{\ast},\beta^{\ast})\;\leq\;-\frac{\varepsilon}{2}\;<\;0.$
This implies that $(\alpha^{\ast},\beta^{\ast})\in\mathcal{A}(\phi)$ for all $\phi\in[0,\phi_{0})$, so $\mathcal{A}(\phi)$ is non-empty in a neighbourhood of zero.
(iii) As $\phi$ increases, $\mathcal{A}(\phi)$ shrinks and is empty for $\phi$ close to one under the stated condition. Suppose that at $\phi=1$ the public signals generated under $E$ and $S$ are observationally equivalent for every $\theta$, in the sense that they induce the same distribution over public histories conditional on $\theta$. In this case the continuation value under $S$ and under $E$ is the same at any public state: given identical signal processes, the law of motion for public beliefs and the worker’s equilibrium continuation value function coincide across the two actions. Hence at $\phi=1$ the sign of $\Delta^{1}(\theta;\alpha,\beta)$ is determined entirely by the current-period payoff difference between $S$ and $E$.
Let $w^{1}(\alpha,\beta)$ denote the equilibrium wage under $E$ at $(\alpha,\beta)$ when $\phi=1$. Under $S$, type $\theta$ receives a Bernoulli payoff (success $1$, failure 0) with mean $\theta$; under our normalization $u(0)=0$ and $u(1)=1$, the expected current utility under $S$ is $u_{S}(\theta)=\theta u(1)+(1-\theta)u(0)=\theta$. Under $E$, the current utility is $u(w^{1}(\alpha,\beta))$. Under competitive wage posting and a non-degenerate talent distribution, zero-profit conditions imply that $w^{1}(\alpha,\beta)<1$ at any state where the posterior places positive mass on $\theta<1$. It follows that
$\Delta^{1}(1;\alpha,\beta)\;=\;u_{S}(1)-u(w^{1}(\alpha,\beta))\;=\;1-u(w^{1}(\alpha,\beta))\;>\;0$
for all such states: the highest types strictly prefer $S$ to $E$ when the two actions generate the same public signals. Thus there is no public state $(\alpha,\beta)$ at which $E$ is weakly preferred to $S$ for all $\theta\in[0,1]$, so $\mathcal{A}(1)=\emptyset$.
Finally, continuity of $\Delta^{\phi}$ in $\phi$ implies that the correspondence $\phi\mapsto\mathcal{A}(\phi)$ is upper hemicontinuous in the Hausdorff topology: if $\phi_{n}\to\phi$ and $(\alpha_{n},\beta_{n})\in\mathcal{A}(\phi_{n})$ with $(\alpha_{n},\beta_{n})\to(\alpha,\beta)$, then $\Delta^{\phi}(\theta;\alpha,\beta)\leq 0$ for all $\theta$. Since $\mathcal{A}(1)=\emptyset$, upper hemicontinuity implies that there exists $\phi_{1}<1$ such that $\mathcal{A}(\phi)=\emptyset$ for all $\phi\in(\phi_{1},1]$. Moreover, as $\phi$ increases, the informativeness of the in-firm signal in the sense of Blackwell increases and the continuation values under $E$ and $S$ converge, so the set of public states at which $E$ strictly dominates $S$ for all types can only shrink. This yields the claimed monotonicity and completes the proof. ∎
Appendix B Details for truncated Beta and the fixed point
With $f(t)\propto t^{\alpha-1}(1-t)^{\beta-1}$, $F(c)=\int_{0}^{c}f$ and $A(c)=\int_{0}^{c}tf$, we have $m(c)=A/F$ and
$m^{\prime}(c)=\frac{f(c)}{F(c)}\,[c-m(c)]>0,\qquad\frac{d}{dc}\big(u(m(c))\big)=u^{\prime}(m(c))\,m^{\prime}(c)>0.$
These identities imply the monotonicity used in the existence/uniqueness proof and are convenient for comparative statics.
At any $(\alpha,\beta)$, define $\phi(w)$ by $\Delta(\phi(w);\alpha,\beta,w)=0$.
By the implicit function theorem,
$\phi^{\prime}(w)=-\frac{\partial\Delta/\partial w}{\partial\Delta/\partial\theta}<0,$
[…]
Appendix C Comparative statics with worker learning and partial observability
Let the worker hold a private posterior $H_{t}$ with mean $m_{t}$. Replace $\theta$ in (4) with $m_{t}$ in the current term and in the transition probabilities. The mapping $m\mapsto U_{S}(m)-U_{E}(m)$ is strictly increasing by the same argument as in Lemma 1, so a unique cutoff $\hat{m}(\alpha,\beta)$ exists. Since employment freezes the public state $(\alpha,\beta)$, the same comparison repeats next period; local absorption follows.
If $E$ generates a Bernoulli public signal with informativeness $\phi\in[0,1)$, posterior updating at $E$ weakly increases the option value of employment. For sufficiently small $\phi$ (bounded away from 1), the signal is too noisy to overturn absorption: the wage still prices the selected pool and the single-crossing comparison is preserved. As $\phi\uparrow 1$, the absorbing region shrinks and vanishes in the limit where $E$ is as informative as $S$.